### v1

# GDSR: Global-Detail Integration through Dual-Branch Network with Wavelet Losses for Remote Sensing Image Super-Resolution

Abstract
In recent years, deep neural networks, including Convolutional Neural Networks, Transformers, and State Space Models, have achieved significant progress in Remote Sensing Image (RSI) Super-Resolution (SR). However, existing SR methods typically overlook the complementary relationship between global and local dependencies. These methods either focus on capturing local information or prioritize global information, which results in models that are unable to effectively capture both global and local features simultaneously. Moreover, their computational cost becomes prohibitive when applied to large-scale RSIs. To address these challenges, we introduce the novel application of Receptance Weighted Key Value (RWKV) to RSI-SR, which captures long-range dependencies with linear complexity. To simultaneously model global and local features, we propose the Global-Detail dual-branch structure, GDSR, which performs SR reconstruction by paralleling RWKV and convolutional operations to handle large-scale RSIs. Furthermore, we introduce the Global-Detail Reconstruction Module (GDRM) as an intermediary between the two branches to bridge their complementary roles. In addition, we propose Wavelet Loss, a loss function that effectively captures high-frequency detail information in images, thereby enhancing the visual quality of SR, particularly in terms of detail reconstruction. Extensive experiments on several benchmarks, including AID, AID_CDM, RSSRD-QH, and RSSRD-QH_CDM, demonstrate that GSDR outperforms the state-of-the-art Transformer-based method HAT by an average of 0.05 dB in PSNR, while using only 63% of its parameters and 51% of its FLOPs, achieving an inference speed 2.9 times faster. Furthermore, the Wavelet Loss shows excellent generalization across various architectures, providing a novel perspective for RSI-SR enhancement.
Index Terms:
remote sensing image, super-resolution, global-detail reconstruction, dual-branch.I Introduction
The rapid development of remote sensing technologies has resulted in an ever-increasing availability of satellite and aerial images, which play a crucial role in a variety of applications, such as environmental monitoring [1], urban planning [2], and disaster management [3]. However, the resolution of these remote sensing images is often constrained by the limitations of imaging sensors, atmospheric interference, and satellite bandwidth. Consequently, SR techniques, which seek to reconstruct high-resolution (HR) images from low-resolution (LR) observations, have garnered significant interest from both academic and industrial sectors [4, 5]. In particular, remote sensing image super-resolution (RSISR) has emerged as a pivotal research area that enhances spatial resolution and reveals finer details that are essential for accurate analysis.
Traditional SR methods have been dominated by Convolutional Neural Networks (CNNs) [6, 7] due to their strong local feature extraction and efficient computation. Models such as SRResNet [8] and EDSR [9] have demonstrated that deeper CNNs can progressively enhance SR performance. However, they suffer from the Effective Receptive Field (ERF) problem (as depicted in Fig. 1 (a)). Although deeper networks theoretically have larger receptive fields, only a small portion is effectively utilized during feature extraction, which limits their ability to capture long-range dependencies. This becomes particularly challenging in RSIs, where diverse spatial patterns and textures overwhelm the limited modeling capacity of CNN-based methods [10, 11].
In recent years, Transformers have gained significant traction in computer vision [14, 15], including RSISR [10, 16]. They excel at capturing global dependencies by modeling long-range interactions through self-attention mechanisms. However, the quadratic complexity of the self-attention operation can become a bottleneck when processing large-scale satellites or aerial images. To address this, many Transformer-based models aim to reduce computational demands by limiting attention to smaller windows [14], which in turn reduces the effective receptive field (as shown in Fig. 1 (b)) and conflicts with the goal of capturing long-range dependencies. Additionally, HAT [15] combines channel attention and window-based self-attention with an overlapping cross-attention module to capture long-range dependencies. However, although transformers can capture the global context, they struggle with computational complexity and may lose structural information during sequence transformation [17], necessitating a re-evaluation of existing methods and the development of novel approaches.
The linear architecture of Transformers like Mamba-based models [18] and RWKV-based models [13] serves as an efficient alternative to Transformers. Mamba-based models leverage State Space Models (SSMs) [19] to causally capture long-range dependencies with linear computational complexity, with recent studies suggesting that they can achieve performance comparable to or even better than Transformer-based models [18, 20]. However, their inherent sequence modeling in SSMs still poses challenges in achieving an optimal ERF for 2D images (as depicted in Fig. 1 (c)). RWKV-based models [21, 22], originally developed for natural language processing (NLP), has emerged as a potential alternative to the transformer and amba. RWKV improves efficiency with linear complexity for long-range dependencies and better captures local dependencies, enhancing performance on large-scale NLP and vision tasks. Recent studies have shown that RWKV-based models outperform Transformer-based models and linearly scaled Mamba-based models in terms of effectiveness and efficiency [22, 23]. Despite pioneering efforts to adopt RWKV for vision tasks, the adaptation of RWKV to the field of RSISR remains unexplored.
While CNNs, Transformers, Mamba, and RWKV networks each offer distinct advantages, the challenges of effectively managing receptive fields and addressing the ERF limitations remain open research questions in RSISR. Notably, a recent study [24] has shown that multi-head self-attention (MSA) and convolutional layers exhibit complementary filtering behaviors. Specifically, MSAs act as low-pass filters, emphasizing broader structures and smoothing noise, whereas conv layers serve as high-pass filters, focusing on fine-grained textures and edges. These contrasting properties make the combination of MSAs and conv layers particularly beneficial for improving model performance. Moreover, high-frequency details are crucial for accurately reconstructing remote sensing images, as they play a key role in preserving fine-grained information. However, existing approaches struggle to simultaneously model long-range dependencies while effectively capturing high-frequency details. Therefore, a natural question arises: Can a more efficient yet effective solution be developed to capture long-range dependencies across large-scale RSIs with convolution, which directly handles 2D spatial structures and offers computational advantages, while also efficiently reconstructing high-frequency details?
In this paper, we propose a novel approach that integrates global representations and detailed extraction with wavelet losses to enhance the SR performance for remote sensing imagery. Our main contributions can be summarized as follows:
1)
The RWKV is introduced for remote sensing image super-resolution, along with GDSR, a novel architecture designed to effectively combine global and local information through a dual-branch structure.
2)
A Global–Detail Reconstruction Module (GDRM) is proposed to efficiently integrate information from the dual branches. By merging complementary features, this module enhances the synergy between the branches, thereby significantly improving the overall performance of RSISR tasks.
3)
Wavelet loss is introduced as a simple yet powerful loss function for RSISR. This approach effectively enhanced both the PSNR and visual quality, and demonstrated robust performance improvements across various network architectures. Moreover, its ease of implementation and adaptability make it a practical and highly effective solution for advancing RSISR techniques.
We demonstrate the empirical validity of GDSR by comparing it with various competitive CNN-, GAN-, Transformer-, Mamba-, and hybrid-based models on two RSI benchmark datasets using two degradation methods showing SOTA results. In addition, we conducted extensive ablation experiments (as detailed in Section IV-C) to investigate the approach of global-detail integration and its effect on RSI-SR, and to examine the effect of Wavelet Loss on various model architectures. We believe our work will advance efficient large-scale RSI-SR, improving both fidelity and processing speed, while ensuring robustness and practical applicability for real-world remote sensing tasks.
II Related Work
II-A Remote Sensing Image Super-resolution
Recent advances in deep learning have greatly improved RSI-SR, with methods typically falling into CNN-based, Transformer-based, and Mamba-based categories. CNN-based methods, inspired by SRCNN [6], have used modules like residual connections [7], dense structures [25], and attention mechanisms [26], [27] to enhance performance. However, they are limited by their inability to capture long-range dependencies, which hinders their effectiveness in large-scale RSI data. Transformer-based methods, leveraging self-attention (SA) [28], have shown promise in modeling global dependencies, with approaches like window-based SA [15] and recursive SA [29] attempting to address the quadratic complexity. However, SA remains inefficient for high-resolution images, often sacrificing global modeling capability. To overcome these limitations, Mamba-based methods, which offer linear complexity for global modeling, present a compelling solution. MambaIR [18] successfully introduced Mamba into the SR domain, while FreMamba [20] pioneers the exploration of Mamba’s potential for RSI-SR, extending it with frequency analysis. Due to the unidirectional nature of SSM, Mamba’s global dependency modeling still faces challenges in achieving accurate results for RSI-SR.
II-B RWKV in computer vision
Recent studies [22] indicate that RWKV can outperform both Transformers and Mamba in NLP tasks. Vision-RWKV [30] has been successfully applied to vision tasks, showing superior performance compared to vision Transformers with lower computational complexity. Several RWKV-based models, such as Diffusion-RWKV [31] for image generation and Restore-RWKV [13] for medical image restoration, have demonstrated promising results. However, the potential of RWKV for RSI-SR remains unexplored. Given that RSI exhibits diverse and complex characteristics, with texture information often subtler than in natural images, it presents unique challenges for SR tasks. Moreover, critical high-frequency details tend to be lost in deep learning models, emphasizing the need for a more effective approach to capture these crucial components. This paper pioneers the exploration of RWKV’s potential for RSI-SR by introducing a novel dual-branch architecture that combines RWKV with convolutional operations.
II-C Wavelet Transforms in Deep Learning
The wavelet transform (WT) [32] has been a fundamental tool for signal processing since the 1980s and has more recently been integrated into neural network architectures for various tasks. Several studies have explored wavelet-domain methods for image restoration and SR tasks, recognizing the need to treat low- and high-frequency features differently. Methods like Wavelet-SRNet [33], WDST [34] and WDRN [35] predict wavelet coefficients to enhance SR performance. More recently, WGSR [36] introduced wavelet subband losses in GAN-based SR models to better control artifacts. In addition, WTConv [37], a novel layer leveraging the wavelet transform to effectively expand the receptive field of convolutions, has been proposed as a drop-in replacement for depth-wise convolutions in CNNs, improving performance across several key computer vision tasks. In this paper, we propose a novel Wavelet Loss that directly optimizes the reconstruction of wavelet-domain features. Our approach is designed to complement SR models based on various architectures, including CNN, GAN, Transformer, and Mamba, as well as their hybrid combinations, which have demonstrated superior performance in RSISR. Unlike previous methods, our Wavelet Loss seamlessly integrated wavelet-based supervision into the end-to-end SR training framework to ensure effective high-frequency detail recovery.
III Methodology
In this section, we first introduce the overall structure of the proposed GDSR and then described three important modules in GDSR, namely RGEG, RDEG, and GDRM. Finally, we detail our designed loss function.
III-A Overview of GDSR
As illustrated in Fig. 2, GDSR begins by extracting the shallow feature from the LR input using a convolutional layer. It then derives the deep features and through parallel branches: global extraction branch and details extraction branch, respectively. These features are subsequently fused using the GDRM. A global residual connection is then applied to combine the low-level and deep features, and finally, the pixel-shuffle method is employed for upsampling to generate the SR output.
The SR process of GDSR can be mathematically described as:
where denotes the upsampling function and denotes the global-detail reconstruction operation. is the reconstructed SR image with a scaling factor of .
III-B Residual Details Extraction Group
Deep convolutional networks have achieved remarkable advancements in SR tasks, with residual structures proving effective in addressing network degradation issues [9]. To harness these advantages, we propose RDEGs, which are structured as sequences of residual units, each referred to as a Residual Convolution Block (RCB). RCBs are specifically designed to enhance the recovery of fine details from LR inputs while maintaining stable and efficient model training.
A single RCB can be represented as:
where and represent the input and output feature maps of the -th RCB, respectively. denotes a learnable scale factor that controls the amount of information passed through the residual connection. and denote the convolution operations in the RCB and refers to the Leaky ReLU activation function.
We concatenate RCBs to form an RDEG module. The RDEG can be mathematically expressed as:
where . Here, represents the feature map of the -th RDEG, denotes the output of the -th RCB within the -th RDEG.
III-C Residual Global Extraction Group
Inspired by the recent success of RWKV in computer vision applications [30, 13], we seek to utilize its strengths in modeling long-range dependencies in RSIs. As illustrated in Fig. 2, the proposed RGEG consists of multiple Residual RWKV Blocks (R-RWKVBs). The spatial mix (sm) module is designed to establish long-range dependencies among tokens across the spatial dimension. Given an input feature flattened into a one-dimensional sequence , where represents the total number of tokens, the spatial mix module begins by applying layer normalization (LN) followed by the Omni-Shift operation [13]. The output is then passed through three parallel linear projection layers to obtain the receptance , key , and value :
where , , and are linear projection matrices. The key and value are then processed through the Re-WKV attention mechanism [13] to compute the global attention output . Finally, the receptance , after gating through a sigmoid function , modulates the attention output via element-wise multiplication:
where represents the spatial mix output, denotes element-wise multiplication, and is the output projection matrix. The channel mix (cm) module focuses on channel-wise feature fusion. Similar to the spatial mix module, it starts with LN and Omni-Shift, and the receptance , key , and value are computed as follows:
where , , and are linear projection matrices, and is the squared ReLU activation function, which enhances nonlinearity. Notably, the transformation from to to involves a multi-layer perception (MLP) consisting of , , and , facilitating channel-wise feature fusion. The final channel mix output is obtained through:
where is the output projection matrix. Finally, RGEG can be formulated as
where and denote the feature map of the -th RGEG and -th RGEG, respectively. is the -th R-RWKVB. is a convolutional layer that serves to enhance the translational equivariance of the RWKV layer, while and are learnable scale factors.
III-D Global–Detail Reconstruction Module
Remote sensing images often exhibit intricate spatial details and large-scale contextual information, necessitating a specialized module to harmonize these features for enhanced reconstruction performance. To effectively integrate the complementary features extracted by the RDEG and the RGEG, the GDRM is designed as a fusion mechanism tailored for RSI-SR tasks.
To harmonize and align these two feature representations, the GDRM employs a Permuted Spatial Attention Module (PSAM), which is designed to be computationally efficient and structurally simple. As illustrated in Fig. 3, the module begins by applying a permute operation to the input feature map, reorganizing it into three distinct formats, each capturing specific spatial relationships: Height-Width (HW), Channel-Width (CW), and Height-Channel (HC). Spatial Attention [38] is subsequently applied to each permuted representation independently, enabling the modeling of dependencies along these dimensions. The processed features are then permuted back to their original format, summed together, and scaled by a coefficient. Formally, the aligned feature maps are computed as:
{align}
[…]
where denotes the spatial attention operation and is a fixed coefficient. As illustrated in Fig. 4, the GDRM receives two input feature maps: one capturing local details from the RDEG branch and another representing global contextual information from the RGEG branch. The GDRM then processes these feature maps through a series of operations, which can be expressed as:
[…]
III-E Wavelet Loss
Due to the unique characteristics of remote sensing imagery, preserving fine-grained details across multiple scales is critical for enhancing spatial resolution while preserving fine-grained spatial details and structural consistency. To address this challenge, we introduce an integrable wavelet-based loss function designed for versatile and effective application across diverse SR models, leveraging the Haar wavelet transform to capture high-frequency details essential for visually and quantitatively improved RSI-SR results.
As illustrated in Fig. 5, the SWT generates one low-frequency (LF) subband, LL, representing the coarse structure of the image, and three high-frequency (HF) subbands, LH, HL, and HH, which encode horizontal, vertical, and diagonal details, respectively. As illustrated in Fig. 6, the wavelet loss function is formulated as the distance between the corresponding Haar wavelet subbands of the generated SR image and the HR image. The total wavelet loss is expressed as:
where and are the HR and SR images, denotes the Haar wavelet transform, indexes the subbands, and are scaling factors that balance the importance of each subband. The total wavelet loss is averaged over a minibatch size represented by . By emphasizing high-frequency subbands, the loss effectively enhances edge and texture preservation, while the low-frequency subband ensures structural consistency. The overall loss for the training with wavelet loss is given by
IV Experiment
County | Longitude (°E) | Latitude (°N) |
Maqin | 100.219 to 100.285 | 34.264 to 34.7 |
Gande | 100.047 to 100.719 | 34.054 to 34.204 |
Dari | 98.7845 to 98.996 | 33.485 to 33.925 |
Maduo | 97.929 to 98.129 | 34.668 to 34.698 |
Henan | 101.374 to 101.759 | 34.307 to 34.707 |
Gangcha | 100.406 to 100.509 | 37.351 to 37.671 |
Tianjun | 98.409 to 98.624 | 37.45 to 37.4855 |
IV-A Datasets and Evaluation
This paper reports the performance of SR on two remote sensing image datasets, including the publicly available AID dataset [39] and the custom RSSRD-QH dataset.
The AID dataset consists of 30 categories. For each category, the data was randomly split into two equal parts, one for training and the other for testing. Additionally, 20% of the training data was further separated and used as a validation set. Specifically, the training set contained 4,000 images, the validation set 1,000 images, and the test set 5,000 images.
The RSSRD-QH dataset is located in representative watershed units within the Sanjiangyuan region and the Qinghai Lake surrounding area. The geographical distribution of the counties covered by the dataset, including their longitude and latitude ranges, is summarized in Table I. The dataset was collected using unmanned aerial vehicles, with flight altitudes carefully controlled between 30 and 50 meters. The average spatial resolution of the captured imagery is 0.01 meters, providing comprehensive coverage of various landscape elements across different elevation gradients, including terraces, floodplains, gentle slopes, and steep slopes within these watershed units. From each Watershed, we randomly selected 50% of the images for the training set and 50% for the test set, with 20% of the training set randomly selected for validation. This resulted in a training set of 2,242 images, a validation set of 561 images, and a test set of 2,804 images. The HR images in both datasets were 600 × 600 pixels, and the corresponding LR images were 200 × 200 pixels.
The proposed method, along with other competing methods, was evaluated on the test datasets using two classic full-reference metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [40]. Additionally, the LPIPS [41] metric was used to capture the perceptual quality of the images. Notably, PSNR and SSIM were calculated based on the luminance channel (Y) of the YCbCr color space, while LPIPS evaluated the perceptual similarity between images by extracting features from a pre-trained deep network, specifically the AlexNet model [42], which was employed in this study.
IV-B Degradation Model and Implementation Details
We employed both bicubic downsampling and a comprehensive degradation model (CDM) to simulate LR images during training. Based on existing blind SR methods [43, 44] and the characteristics of remote sensing sensors [45], the CDM for LR image synthesis included isotropic Gaussian blur, motion blur, scaling with different interpolation methods, additive Gaussian noise, and JPEG compression noise, all applied in random order.
Our GDSR performed deep feature exploration through 4 RGEGs, RDEGs, and GDRMs. Each RGEG consisted of 6 R-RWKVBs and each RDEG contains 12 RCBs. Empirically, we set the internal channel dimension to = 96. For the GDSR_TC and GDSR_MC, the RGEG branches correspond to the RTSB and ResidualGroup of SwinIR [14] and MambaIR [18], respectively, with all other settings remaining consistent with GDSR.
During training, each HR image was randomly cropped to a size of 192 × 192 pixels, with the corresponding LR image size being 64 × 64 pixels. The batch size was set to 16, and the Adam optimizer was used with = 0.9 and = 0.999. The initial learning rate was set to , which was halved at the 100th epoch, with a total of 200 training epochs. All SR models were implemented in the PyTorch framework. The FLOPs results were calculated using an input tensor of size 1 × 3 × 240 × 240, and the computations were performed using torch-operation-counter [46]. The inference times were tested on 300 random images with a size of 3 × 240 × 240. All experiments were conducted on an NVIDIA A800 80GB GPU.
RCBs | PSNR/SSIM | LPIPS | #Param. | FLOPs | FPS |
1 | 28.96/0.7658 | 0.3514 | 5.86M | 340.25G | 7.3 |
4 | 28.96/0.7657 | 0.3518 | 7.85M | 455.31G | 7.0 |
8 | 29.00/0.7677 | 0.3499 | 10.51M | 608.73G | 6.7 |
12 | 29.03/0.7686 | 0.3476 | 13.17M | 762.14G | 6.4 |
16 | 29.01/0.7680 | 0.3480 | 15.83M | 915.55G | 6.1 |
20 | 29.03/0.7688 | 0.3477 | 18.48M | 1068.97G | 5.8 |
IV-C Ablation Study
In this section, we discuss the proposed GDSR in depth by investigating the effect of its major components and their variants.
IV-C1 Effect of Key Modules
(a) Effect of RDEG. To investigate the effect of RDEG on the model performance, we conducted ablation experiments by varying the number of RCBs in RDEG.
The experimental results, presented in Tables II and III, indicated that the model achieved a PSNR of 28.94 dB and an SSIM of 0.7652 when only the RGEG branch was utilized without incorporating RDEG. Adding a single RCB significantly improved the model’s performance, increasing the PSNR to 28.96 dB and the SSIM to 0.7658. This improvement highlighted the RCB’s capability to capture critical information, thereby enhancing reconstruction quality.
As the number of RCBs increased, further improvements in PSNR and SSIM were observed. For instance, increasing the RCB count from 1 to 20 raised the PSNR from 28.96 dB to 29.03 dB and the SSIM from 0.7658 to 0.7688. However, the rate of improvement diminished as the RCB count increased, and beyond a certain threshold, performance even began to decline. This trend suggested that while adding more RCBs initially benefited the model’s output, the marginal gains decreased and eventually plateaued. Additionally, increasing the number of RCBs led to a linear rise in computational complexity, as evidenced by the increase in parameters, FLOPs, and a decrease in FPS. This highlighted the trade-off between performance gains and computational efficiency. Therefore, determining the optimal number of RCBs is crucial in practical applications, especially in scenarios where computational efficiency is a critical factor.
(b) Effect of GDRM. To evaluate the effectiveness of the GDRM, we conducted ablation experiments separately using the RDEG and RGEG branches. The experimental results were presented in Table III. The comparisons demonstrated that introducing the GDRM significantly improved performance compared to single-branch models. The GDSR model achieved superior PSNR, SSIM, and LPIPS values, indicating that GDSR effectively integrated the local detail information from RDEG and the global contextual information from RGEG. Additionally, the incorporation of PSAM enabled GDRM to fuse the information from both branches more effectively.
(c) Visualization of GDRM. Fig. 7 visualized the features of the RDEG, RGEG branches, and GDRM. We observed that the feature maps from RDEG and RGEG were complementary, while GDRM combined features from both branches. This demonstrated that, within the dual-branch structure of GDSR, each branch focused on different features, and fusing these features enabled effective image reconstruction.
Method | PSNR (dB) | SSIM | LPIPS |
RDEG Branch | 28.94 | 0.7656 | 0.3554 |
RGEG Branch | 28.94 | 0.7652 | 0.3519 |
w/o PSAM | 29.01 | 0.7677 | 0.3485 |
F2B | 29.01 | 0.7681 | 0.3476 |
F2D(GDSR) | 29.03 | 0.7686 | 0.3476 |
Depth | PSNR/SSIM | LPIPS | #Param. | FLOPs | FPS |
1 | 28.89/0.7633 | 0.3565 | 3.65M | 211.64G | 24.6 |
2 | 28.97/0.7663 | 0.3514 | 6.82M | 395.14G | 12.6 |
4 | 29.03/0.7686 | 0.3476 | 13.17M | 762.14G | 6.4 |
6 | 29.03/0.7686 | 0.3463 | 19.52M | 1129.14G | 4.3 |
8 | 29.04/0.7692 | 0.3456 | 25.86M | 1496.14G | 3.2 |
10 | 29.05/0.7695 | 0.3450 | 32.21M | 1863.13G | 2.5 |
Methods | Venue | AID [39] | AID_CDM [39] | RSSRD-QH | RSSRD-QH_CDM | ||||||||
PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | ||
Bicubic | - | 27.89 | 0.7238 | 0.4400 | 23.15 | 0.5238 | 0.6911 | 26.45 | 0.6447 | 0.4243 | 22.34 | 0.3525 | 0.6993 |
SRCNN [6] | CVPR’2014 | 28.22 | 0.7363 | 0.3973 | 23.38 | 0.5455 | 0.6569 | 26.54 | 0.6512 | 0.4241 | 22.44 | 0.3588 | 0.7523 |
SRGAN [8] | CVPR’2017 | 28.80 | 0.7597 | 0.3601 | 23.78 | 0.5648 | 0.6314 | 26.68 | 0.6586 | 0.4170 | 22.64 | 0.3759 | 0.7421 |
EDSR [9] | CVPR’2017 | 28.94 | 0.7653 | 0.3513 | 23.91 | 0.5698 | 0.6139 | 26.70 | 0.6598 | 0.4141 | 22.74 | 0.3817 | 0.7283 |
SwinIR [14] | CVPR’2021 | 28.95 | 0.7658 | 0.3512 | 23.89 | 0.5705 | 0.6169 | 26.70 | 0.6597 | 0.4133 | 22.72 | 0.3780 | 0.7432 |
HAT [15] | CVPR’2023 | 28.99 | 0.7671 | 0.3479 | 23.99 | 0.5737 | 0.6080 | 26.71 | 0.6602 | 0.4111 | 22.79 | 0.3823 | 0.7312 |
MambaIR [18] | ECCV’2024 | 29.02 | 0.7684 | 0.3487 | 24.07 | 0.5763 | 0.6000 | 26.72 | 0.6607 | 0.4124 | 22.84 | 0.3835 | 0.7387 |
ConvFormerSR [16] | TGRS’2024 | 28.90 | 0.7637 | 0.3516 | 23.97 | 0.5710 | 0.6118 | 26.69 | 0.6603 | 0.4090 | 22.77 | 0.3779 | 0.7365 |
FreMamba [20] | TMM’2024 | 28.96 | 0.7672 | 0.3410 | 23.89 | 0.5692 | 0.6148 | 26.70 | 0.6601 | 0.4112 | 22.60 | 0.3660 | 0.7242 |
GDSR_TC (Ours) | - | 28.96 | 0.7660 | 0.3500 | 24.01 | 0.5717 | 0.6126 | 26.71 | 0.6604 | 0.4144 | 22.79 | 0.3776 | 0.7401 |
GDSR_MC (Ours) | - | 28.97 | 0.7668 | 0.3472 | 24.03 | 0.5734 | 0.6080 | 26.71 | 0.6602 | 0.4107 | 22.84 | 0.3819 | 0.7327 |
GDSR (Ours) | - | 29.03 | 0.7686 | 0.3476 | 24.10 | 0.5770 | 0.5986 | 26.72 | 0.6612 | 0.4106 | 22.86 | 0.3861 | 0.7268 |
GDSR_L (Ours) | - | 29.04 | 0.7692 | 0.3456 | 24.15 | 0.5795 | 0.5936 | 26.73 | 0.6615 | 0.4098 | 22.89 | 0.3875 | 0.7253 |
Methods | Venue | AID [39] | AID_CDM [39] | RSSRD-QH | RSSRD-QH_CDM | ||||||||
PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | ||
SRCNN [6] | CVPR’2014 | 28.25 | 0.7369 | 0.3998 | 23.38 | 0.5456 | 0.6609 | 26.55 | 0.6498 | 0.4221 | 22.44 | 0.3590 | 0.7485 |
SRGAN [8] | CVPR’2017 | 28.82 | 0.7605 | 0.3672 | 23.80 | 0.5652 | 0.6458 | 26.69 | 0.6581 | 0.4183 | 22.65 | 0.3760 | 0.7456 |
EDSR [9] | CVPR’2017 | 28.96 | 0.7658 | 0.3524 | 23.94 | 0.5711 | 0.6233 | 26.70 | 0.6596 | 0.4128 | 22.76 | 0.3821 | 0.7357 |
SwinIR [14] | CVPR’2021 | 28.98 | 0.7668 | 0.3533 | 23.90 | 0.5704 | 0.6301 | 26.71 | 0.6592 | 0.4147 | 22.73 | 0.3781 | 0.7488 |
HAT [15] | CVPR’2023 | 29.04 | 0.7692 | 0.3483 | 24.01 | 0.5736 | 0.6183 | 26.72 | 0.6597 | 0.4137 | 22.81 | 0.3829 | 0.7348 |
MambaIR [18] | ECCV’2024 | 29.04 | 0.7690 | 0.3521 | 24.08 | 0.5767 | 0.6114 | 26.72 | 0.6598 | 0.4134 | 22.85 | 0.3842 | 0.7388 |
ConvFormerSR [16] | TGRS’2024 | 28.93 | 0.7647 | 0.3524 | 23.99 | 0.5707 | 0.6281 | 26.71 | 0.6591 | 0.4141 | 22.77 | 0.3770 | 0.7427 |
FreMamba [20] | TMM’2024 | 28.94 | 0.7673 | 0.3397 | 23.89 | 0.5682 | 0.6318 | 26.70 | 0.6599 | 0.4130 | 22.64 | 0.3698 | 0.7130 |
GDSR_TC (Ours) | - | 28.99 | 0.7670 | 0.3522 | 24.02 | 0.5719 | 0.6237 | 26.72 | 0.6595 | 0.4162 | 22.79 | 0.3776 | 0.7437 |
GDSR_MC (Ours) | - | 29.00 | 0.7675 | 0.3512 | 24.06 | 0.5742 | 0.6159 | 26.72 | 0.6591 | 0.4122 | 22.84 | 0.3815 | 0.7395 |
GDSR (Ours) | - | 29.04 | 0.7689 | 0.3522 | 24.12 | 0.5774 | 0.6121 | 26.73 | 0.6603 | 0.4116 | 22.88 | 0.3862 | 0.7313 |
GDSR_L (Ours) | - | 29.06 | 0.7694 | 0.3499 | 24.16 | 0.5791 | 0.6069 | 26.73 | 0.6603 | 0.4119 | 22.90 | 0.3875 | 0.7314 |
Fig. 8 provided visual comparisons under different branches and reconstruction strategies. It was evident that the RGEG branch had a significantly larger receptive field than the RDEG branch and activated more widely distributed pixels for image reconstruction. As a result, the RGEG branch achieved better PSNR and SSIM values. Two Global-Detail reconstruction strategies, F2B and F2D, yielded different results. F2B, which sent the reconstructed features back to both RGEG and RDEG branches, largely retained the wide receptive field benefits of the RGEG branch and similar DI values. However, its SR results were comparable to those of the single RGEG branch. In contrast, F2D, which sent the reconstructed features only back to the RDEG branch, achieved DI values between those of the RGEG and RDEG branches. The SR results of the F2D strategy showed that it correctly reconstructed textures and delivered more desirable outcomes. This finding aligned with previous studies, which suggested that while expanding the model’s receptive field allowed consideration of a broader range of pixels, it could also lead to incorrect texture reconstruction. This indicated that simply leveraging information from more pixels was insufficient for accurate reconstruction; effectively utilizing pixel information within the receptive field was crucial. As shown in the LAM Attribution map in Fig. 8, the F2D strategy, compared to F2B and the RGEG branch, demonstrated a greater contribution of pixels within the corresponding regions to the SR results. In contrast, the distant pixels captured by the wide receptive field of the RGEG branch (e.g., pixels in the upper part of the figure) contributed less to the SR results. This suggested that F2D, by routing features exclusively back to the RDEG branch, enabled the model to utilize information more efficiently and reconstruct accurate textures by correctly leveraging the pixels within the receptive field.
(d) Effect of Network Depth on Performance. To investigate the effect of network depth on the performance of GDSR, we varied the shared depth of both the RDEG and RGEG branches while keeping other design parameters constant. As illustrated in Table IV, the experimental outcomes consistently demonstrated an enhancement in performance with deeper networks. Moreover, an evaluation of the trade-offs in terms of computational complexity was conducted. As the network depth increased, the model’s parameter count, FLOPS, and FPS were also scaled accordingly. These results supported the effectiveness of GDSR’s dual-branch design, where the shared depth allowed for both high-quality super-resolution and manageable computational overhead.
Categoties | Bicubic | SRGAN [8] | EDSR [9] | HAT [15] | MambaIR [18] | ConvFormerSR [16] | FreMamba [20] | GDSR (Ours) | ||||||||
PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |
Airport | 27.43 | 0.7387 | 28.48 | 0.7765 | 28.63 | 0.7823 | 28.69 | 0.7837 | 28.72 | 0.7856 | 28.59 | 0.7805 | 28.68 | 0.7851 | 28.72 | 0.7853 |
Bare Land | 34.45 | 0.8344 | 34.84 | 0.8422 | 34.92 | 0.8439 | 34.94 | 0.8445 | 34.96 | 0.8449 | 34.91 | 0.8436 | 34.93 | 0.8444 | 34.97 | 0.8449 |
Baseball Field | 29.19 | 0.7786 | 30.27 | 0.8094 | 30.42 | 0.8137 | 30.45 | 0.8144 | 30.52 | 0.8160 | 30.35 | 0.8123 | 30.43 | 0.8147 | 30.53 | 0.8160 |
Beach | 31.32 | 0.7947 | 31.86 | 0.8082 | 31.94 | 0.8103 | 31.99 | 0.8117 | 32.00 | 0.8118 | 31.94 | 0.8096 | 31.95 | 0.8113 | 32.00 | 0.8115 |
Bridge | 30.19 | 0.8063 | 31.27 | 0.8312 | 31.46 | 0.8353 | 31.52 | 0.8363 | 31.56 | 0.8372 | 31.42 | 0.8339 | 31.52 | 0.8368 | 31.57 | 0.8371 |
Center | 25.76 | 0.6980 | 27.18 | 0.7590 | 27.38 | 0.7665 | 27.47 | 0.7694 | 27.51 | 0.7706 | 27.33 | 0.7645 | 27.43 | 0.7692 | 27.51 | 0.7705 |
Church | 23.09 | 0.6084 | 24.41 | 0.6807 | 24.56 | 0.6895 | 24.61 | 0.6920 | 24.65 | 0.6941 | 24.52 | 0.6875 | 24.56 | 0.6919 | 24.67 | 0.6947 |
Commercial | 26.72 | 0.7178 | 27.59 | 0.7582 | 27.73 | 0.7645 | 27.77 | 0.7660 | 27.81 | 0.7679 | 27.68 | 0.7623 | 27.72 | 0.7659 | 27.81 | 0.7681 |
D-Residential | 23.45 | 0.6301 | 24.46 | 0.6922 | 24.60 | 0.7004 | 24.64 | 0.7026 | 24.69 | 0.7051 | 24.56 | 0.6985 | 24.60 | 0.7029 | 24.70 | 0.7056 |
Desert | 37.14 | 0.8869 | 37.54 | 0.8933 | 37.58 | 0.8944 | 37.63 | 0.8951 | 37.62 | 0.8949 | 37.58 | 0.8941 | 37.60 | 0.8946 | 37.62 | 0.8950 |
Farmland | 32.32 | 0.7934 | 33.18 | 0.8166 | 33.33 | 0.8212 | 33.35 | 0.8216 | 33.42 | 0.8236 | 33.30 | 0.8202 | 33.39 | 0.8235 | 33.43 | 0.8240 |
Forest | 27.67 | 0.6411 | 27.84 | 0.6514 | 27.90 | 0.6558 | 27.95 | 0.6605 | 27.95 | 0.6604 | 27.88 | 0.6552 | 27.92 | 0.6598 | 27.97 | 0.6616 |
Industrial | 26.29 | 0.7063 | 27.43 | 0.7566 | 27.62 | 0.7647 | 27.67 | 0.7669 | 27.71 | 0.7689 | 27.56 | 0.7621 | 27.62 | 0.7665 | 27.72 | 0.7690 |
Meadow | 32.07 | 0.7100 | 32.33 | 0.7136 | 32.38 | 0.7157 | 32.40 | 0.7167 | 32.41 | 0.7172 | 32.37 | 0.7154 | 32.39 | 0.7175 | 32.42 | 0.7172 |
M-Residential | 26.32 | 0.6578 | 27.40 | 0.7055 | 27.55 | 0.7124 | 27.58 | 0.7136 | 27.63 | 0.7159 | 27.52 | 0.7113 | 27.55 | 0.7140 | 27.64 | 0.7164 |
Mountain | 28.05 | 0.6978 | 28.41 | 0.7127 | 28.48 | 0.7163 | 28.50 | 0.7184 | 28.50 | 0.7183 | 28.46 | 0.7152 | 28.45 | 0.7171 | 28.50 | 0.7183 |
Park | 27.14 | 0.7060 | 27.77 | 0.7332 | 27.87 | 0.7383 | 27.91 | 0.7406 | 27.93 | 0.7415 | 27.83 | 0.7365 | 27.86 | 0.7395 | 27.94 | 0.7422 |
Parking | 24.11 | 0.7310 | 26.18 | 0.8059 | 26.51 | 0.8159 | 26.68 | 0.8205 | 26.76 | 0.8225 | 26.46 | 0.8145 | 26.70 | 0.8226 | 26.77 | 0.8228 |
Playground | 29.33 | 0.7723 | 30.70 | 0.8132 | 30.94 | 0.8200 | 31.03 | 0.8216 | 31.10 | 0.8235 | 30.88 | 0.8182 | 31.03 | 0.8230 | 31.12 | 0.8237 |
Pond | 28.76 | 0.7559 | 29.38 | 0.7749 | 29.46 | 0.7781 | 29.49 | 0.7790 | 29.51 | 0.7796 | 29.44 | 0.7769 | 29.46 | 0.7788 | 29.51 | 0.7796 |
Port | 25.83 | 0.7727 | 26.94 | 0.8207 | 27.11 | 0.8255 | 27.18 | 0.8273 | 27.21 | 0.8283 | 27.05 | 0.8237 | 27.13 | 0.8267 | 27.21 | 0.8278 |
Railway Station | 27.00 | 0.6965 | 27.87 | 0.7347 | 28.04 | 0.7434 | 28.11 | 0.7456 | 28.13 | 0.7474 | 27.99 | 0.7404 | 28.07 | 0.7461 | 28.14 | 0.7477 |
Resort | 26.78 | 0.7160 | 27.67 | 0.7547 | 27.81 | 0.7601 | 27.86 | 0.7622 | 27.89 | 0.7634 | 27.78 | 0.7588 | 27.82 | 0.7619 | 27.91 | 0.7638 |
River | 29.29 | 0.7201 | 29.74 | 0.7378 | 29.81 | 0.7412 | 29.84 | 0.7426 | 29.85 | 0.7430 | 29.79 | 0.7401 | 29.80 | 0.7422 | 29.86 | 0.7431 |
School | 25.81 | 0.7046 | 26.81 | 0.7503 | 26.96 | 0.7568 | 27.00 | 0.7587 | 27.04 | 0.7604 | 26.91 | 0.7546 | 26.94 | 0.7577 | 27.05 | 0.7607 |
S-Residential | 25.71 | 0.5790 | 26.15 | 0.6001 | 26.22 | 0.6055 | 26.25 | 0.6070 | 26.27 | 0.6079 | 26.21 | 0.6046 | 26.23 | 0.6075 | 26.27 | 0.6077 |
Square | 27.51 | 0.7286 | 28.71 | 0.7782 | 28.90 | 0.7847 | 28.96 | 0.7870 | 28.99 | 0.7881 | 28.85 | 0.7831 | 28.92 | 0.7868 | 29.01 | 0.7882 |
Stadium | 26.23 | 0.7241 | 27.55 | 0.7771 | 27.75 | 0.7845 | 27.82 | 0.7873 | 27.85 | 0.7882 | 27.68 | 0.7823 | 27.78 | 0.7867 | 27.85 | 0.7881 |
Storage Tanks | 25.15 | 0.6667 | 26.25 | 0.7190 | 26.39 | 0.7255 | 26.41 | 0.7264 | 26.45 | 0.7287 | 26.33 | 0.7233 | 26.39 | 0.7274 | 26.45 | 0.7286 |
Viaduct | 26.77 | 0.6817 | 27.70 | 0.7225 | 27.88 | 0.7310 | 27.91 | 0.7325 | 27.98 | 0.7361 | 27.81 | 0.7277 | 27.87 | 0.7326 | 27.98 | 0.7363 |
Average | 27.89 | 0.7238 | 28.80 | 0.7597 | 28.94 | 0.7653 | 28.99 | 0.7671 | 29.02 | 0.7684 | 28.90 | 0.7637 | 28.96 | 0.7672 | 29.03 | 0.7686 |
IV-C2 Effect of Wavelet Loss
To evaluate the effectiveness of incorporating Wavelet Loss in training, ablation experiments were conducted comparing models trained with Wavelet Loss to those using only Loss. These experiments spanned multiple datasets and degradation models, employing a variety of architectures. The results consistently demonstrated that Wavelet Loss improves PSNR, with SSIM also showing notable enhancement in most cases. On the RSSRD-QH dataset, a decrease in SSIM performance was observed, which was attributed to the complexity of basin images, characterized by intricate textures and diverse structures. Wavelet Loss emphasized high-frequency details, which didn’t align with SSIM’s focus on global structural similarity. A decline in LPIPS was also observed across all experiments with Wavelet Loss, due to its emphasis on high-frequency details and structural accuracy. The prioritization of fine-grained features introduced perceptual differences, as captured by LPIPS. Nonetheless, the substantial improvements in PSNR and SSIM underscored the overall effectiveness of Wavelet Loss in producing high-quality SR images, guiding the correct reconstruction of textures across various architectures.
As shown in Figure 9, despite MambaIR’s global receptive field and its superior performance in experiments, the model still struggled to accurately reconstruct certain textures without Wavelet Loss. The inclusion of Wavelet Loss significantly improved the reconstruction, particularly in the high-frequency components (LH, HL, HH), where fine-grained details were preserved. This highlights Wavelet Loss’s role in guiding the model to recover intricate structures and accurately reconstruct complex textures, thereby enhancing overall image quality.
IV-D Comparisons With State-of-the-Art
IV-D1 Comparative Methods
CNN-based models, including SRCNN [6] and EDSR [9]; GAN-based models, including SRGAN [8]; Transformer-based models, including SwinIR [14] and HAT [15]; CNN-Transformer hybrid models, including ConvFormerSR [16]; and Mamba-based models, including MambaIR [18] and FreMamba [20], were involved in evaluating the SR performance of our GDSR against SOTA methods on RSI-SR. We also report the results of our model variants, GDSR_TC and GDSR_MC, as well as the large-scale variant GDSR_L.
IV-D2 Quantitative Evaluations
The quantitative results on the AID and RSSRD-QH datasets are presented in Table V. Our proposed GDSR and GDSR_L models achieved superior performance across all metrics, demonstrating their remarkable SR capabilities in diverse remote sensing tests. On the AID and RSSRD-QH datasets, GDSR achieved performance comparable to MambaIR. On the AID_CDM and RSSRD-QH_CDM datasets, GDSR’s advantages became more pronounced. It achieved PSNR, SSIM, and LPIPS values of 24.10, 0.5770, and 0.5986, respectively, on AID_CDM, and PSNR and SSIM values of 22.86 and 0.3861, respectively, on RSSRD-QH_CDM, significantly outperforming other methods. GDSR_L consistently led across all tests, and both GDSR_TC and GDSR_MC achieved performance comparable to FreMamba on the AID and RSSRD-QH datasets while demonstrating substantial improvements on the AID_CDM and RSSRD-QH_CDM datasets. Specifically, GDSR_MC achieved a 0.14 dB PSNR improvement on AID_CDM and a 0.24 dB improvement on RSSRD-QH_CDM compared to FreMamba.
To further evaluate the generalization capability of the super-resolution models across diverse remote sensing scenarios, we report the PSNR and SSIM results across 30 scene categories in the AID dataset, as shown in Table VII. As evidenced, GDSR exhibited stronger generalization capability compared to state-of-the-art methods, achieving the best performance in nearly all remote sensing scenarios. Specifically, for the ”Parking” and ”Playground” scenes, GDSR outperformed HAT by 0.09 dB in PSNR and by 0.0023 and 0.0021 in SSIM, respectively. Additionally, in the most complex categories, such as ”Church” and ”DenseResidential,” GDSR also achieved the best performance.
Moreover, as the receptive field of the models increased, performance trended upward for CNN-based EDSR, Transformer-based SwinIR, and Mamba-based MambaIR. This aligned with our experimental results on the RGEG Branch and RDEG Branch, where larger receptive fields enabled the activation of more pixels for image reconstruction, enhancing performance. However, the results from MambaIR and FreMamba, along with the experiments on F2B and F2C, indicated that spatial global modeling in complex remote sensing images had reached a bottleneck. Simply increasing the receptive field no longer yielded significant performance gains. A more practical solution was to effectively utilize large receptive fields to activate the appropriate regions and pixels for accurate image reconstruction. The outstanding generalization capability and superior performance of GDSR aligned with our motivation to address the challenges of large-scale remote sensing images efficiently through the introduction of a global-detail dual-branch structure.
IV-D3 Complexity and Efficiency Evaluation
To evaluate the complexity and computational efficiency of our proposed models, we conducted quantitative assessments across various metrics. As shown in Table VIII, the bicubic interpolation technique inherently avoids the need for parameter optimization, resulting in the fastest inference speed within our evaluation scope. The CNN-based model achieved the highest inference speed among all tested deep learning models, primarily due to the inherent efficiency of convolutional parallel operations. The EDSR model, with its deep residual structure, had the largest number of parameters and FLOPs, yet achieved inference speeds several times faster than other architectures.
Transformer-based models, such as SwinIR and HAT, exhibited slower inference speeds compared to CNN-based models but achieved significant performance gains. Models based on Mamba, including MambaIR and FreMamba, leveraged their linear complexity and optimized inference processes, delivering superior performance along with relatively faster inference speeds. ConvFormerSR, which integrated Swin Transformer with CNN architectures, achieved faster inference speeds while maintaining competitive performance.
GDSR, by incorporating RWKV with CNN, not only reduced the parameter count but also achieved inference speeds several times faster than Transformer- and Mamba-based models, while delivering enhanced performance. GDSR_TC and GDSR_MC further optimized performance by combining Swin Transformer and Mamba with CNN, leading to a reduction in parameter size and improved inference speed, without sacrificing performance. GDSR_L, which increased the network depth, consistently outperformed other models in experimental evaluations while maintaining reasonable inference speeds. As demonstrated in Tables II, III, and IV, these results underscore the superiority of the Global-Detail dual-branch architecture design.
Comparing Tables VIII and V reveals that the complexity of our proposed model is comparable to most state-of-the-art models, while its efficiency is superior, and its performance is significantly better. These results highlight the practical potential of our model, demonstrating an effective balance between computational efficiency and performance, making it well-suited for real-world applications.
Model | #Param. | FLOPs | FPS |
Bicubic | - | - | 21505.6 |
SRCNN [6] | 0.02M | 10.47G | 533.1 |
SRGAN [8] | 1.55M | 107.68G | 151.9 |
EDSR [9] | 43.68M | 2519.95G | 15.5 |
SwinIR [14] | 16.62M | 1005.01G | 2.3 |
HAT [15] | 20.81M | 1467.38G | 2.2 |
MambaIR [18] | 15.11M | 838.91G | 2.8 |
ConvFormerSR [16] | 16.49M | 994.90G | 3.2 |
FreMamba [20] | 11.46M | 516.21G | 2.5 |
GDSR_TC (Ours) | 12.83M | 758.82G | 3.3 |
GDSR_MC (Ours) | 13.99M | 788.28G | 4.1 |
GDSR (Ours) | 13.17M | 762.14G | 6.4 |
GDSR_L (Ours) | 25.86M | 1496.14G | 3.2 |
IV-D4 Qualitative Results
Visual comparisons on AID, AID_CDM, and RSSRD-QH_CDM with a scale factor of ×3 are presented in Fig. 10, Fig. 11, and Fig. 12. From these visual results, it was evident that our proposed GDSR was capable of restoring sharp edges and showcasing richer textures, particularly in capturing critical high-frequency details in remote sensing images. For instance, when comparing the reconstructed ”playground_161” in Fig. 10, the closest competing Transformer-based method HAT, the hybrid CNN-Transformer method ConvFormerSR, and the Mamba-based model MambaIR all failed to reconstruct the high-resolution lines on the runway. Only the proposed GDSR accurately restored these details. Furthermore, as illustrated in Fig. 11 for ”viaduct_398,” under more complex degradation scenarios, GDSR effectively reconstructed the contours, whereas other SR models failed to accurately represent the detailed structure. In Fig. 12, a comparison of the visual results for the RSSRD-QH_CDM dataset showed that the original image, after CDM degradation, appeared heavily blurred and lost its distinctive columnar features. Only the proposed GDSR was able to accurately recover the outline of the original image.
The LAM comparisons in Fig. 13, based on Figs. 10 and 12, demonstrated that GDSR benefitted from the superior long-range modeling capability of RGEG, enabling it to leverage more pixels during the SR process. Additionally, the strong detail extraction capability of RDEG allowed GDSR to utilize pixels correctly for reconstruction. Specifically, for the RSSRD-QH_CDM dataset, when large-scale and global information was present, CNN- and Transformer-based SR networks struggled to extract sufficient information, and while the Mamba-based SR network activated a large number of pixels, they failed to use them correctly for reconstruction. In contrast, thanks to the effective integration of RGEG and RDEG, the GDSR-generated results significantly outperformed other SR models. Moreover, GDSR surpassed MambaIR in terms of a diffusion index by 25.6193, indicating that by incorporating an RWKV with a larger receptive field into RGEG, GDSR could better capture global information in images with more such features, leading to a higher diffusion index. These visual comparisons further validated the effectiveness of GDSR in capturing and reconstructing details in remote sensing images, highlighting the powerful pixel utilization capability brought by our global-detail dual-branch structure.
IV-D5 Real-World Image Testing on the Sanjiangyuan Region
To further evaluate the model’s performance in real-world scenarios, we tested it on remote sensing images from real-world environments. As shown in Fig. 14, the image depicts an alpine meadow grassland type, with surface features such as cracks, scattered rocks, and numerous small ”mouse holes.” These mouse holes are depressions or openings on the ground, and the distribution of rodents is significant for studying the primary grassland type, the alpine meadow, in Qinghai Province. Their presence can impact the biodiversity, productivity, and ecological function of the alpine meadows. However, due to the limitations in drone flight conditions and resolution, the raw images suffer from motion blur and loss of details. These issues hinder accurate observation of key features such as surface textures and the depth of the mouse holes. By applying SR techniques, our GDSR effectively enhanced the clarity of these fine details, sharpened the edges around the mouse holes, and restored the surface texture. This improvement is crucial for accurate environmental analysis, as the restored landscape provides more precise data for studying biodiversity conservation and the enhancement of ecological functions in the alpine meadow ecosystem of Qinghai Province.
V Conclusion
In this study, we first introduce the Receptance Weighted Key Value for remote sensing image super-resolution. GDSR effectively models the global dependencies of large-scale remote sensing images while incorporating local details, all with linear complexity and significantly improved inference speed. Specifically, we design an efficient and effective Global-Detail dual-branch architecture that correctly activates more pixels spatially for image reconstruction. To combine global and local representations, we propose GDRM, which adaptively adjusts the features of the dual branches and fuses them. Additionally, we introduce Wavelet Loss to help the model better capture high-frequency details in remote sensing image super-resolution tasks. Extensive quantitative and qualitative experiments on the AID, AID_CDM, RSSRD-QH, and RSSRD-QH_CDM benchmarks demonstrate that our GDSR outperforms state-of-the-art CNN-, Transformer-, and Mamba-based SR models in remote sensing image super-resolution tasks. The proposed Wavelet Loss also enhances the performance of various models in remote sensing image super-resolution tasks.
This study focuses solely on ×3 SR, limiting the flexibility to explore SR at different scales. Experiments show that excessive high-frequency components introduced by Wavelet Loss can lead to a decline in perceptual quality. In future work, we plan to extend our GDSR to include more scaling factors, further demonstrating its robustness and effectiveness. Additionally, we aim to optimize Wavelet Loss to improve its perceptual enhancement.

### v4

@@
 To address these challenges, we introduce the novel application of Receptance Weighted Key Value (RWKV) to RSI-SR, which captures long-range dependencies with linear complexity.
-To simultaneously model global and local features, we propose the Global-Detail dual-branch structure, GDSR, which performs SR reconstruction by paralleling RWKV and convolutional operations to handle large-scale RSIs.
+To simultaneously model global and local features, we propose the Global-Detail dual-branch structure, GDSR, which performs SR by paralleling RWKV and convolutional operations to handle large-scale RSIs.
 Furthermore, we introduce the Global-Detail Reconstruction Module (GDRM) as an intermediary between the two branches to bridge their complementary roles.
-In addition, we propose Wavelet Loss, a loss function that effectively captures high-frequency detail information in images, thereby enhancing the visual quality of SR, particularly in terms of detail reconstruction.
-Extensive experiments on several benchmarks, including AID, AID_CDM, RSSRD-QH, and RSSRD-QH_CDM, demonstrate that GSDR outperforms the state-of-the-art Transformer-based method HAT by an average of 0.05 dB in PSNR, while using only 63% of its parameters and 51% of its FLOPs, achieving an inference speed 2.9 times faster.
-Furthermore, the Wavelet Loss shows excellent generalization across various architectures, providing a novel perspective for RSI-SR enhancement.
+In addition, we propose the Dual-Group Multi-Scale Wavelet Loss, a wavelet-domain constraint mechanism via dual-group subband strategy and cross-resolution frequency alignment for enhanced reconstruction fidelity in RSI-SR.
+Extensive experiments under two degradation methods on several benchmarks, including AID, UCMerced, and RSSRD-QH, demonstrate that GSDR outperforms the state-of-the-art Transformer-based method HAT by an average of 0.09 dB in PSNR, while using only 63% of its parameters and 51% of its FLOPs, achieving an inference speed 3.2 times faster.
 Index Terms:
@@
 Consequently, SR techniques, which seek to reconstruct high-resolution (HR) images from low-resolution (LR) observations, have garnered significant interest from both academic and industrial sectors [4, 5].
-In particular, remote sensing image super-resolution (RSISR) has emerged as a pivotal research area that enhances spatial resolution and reveals finer details that are essential for accurate analysis.
-Traditional SR methods have been dominated by Convolutional Neural Networks (CNNs) [6, 7] due to their strong local feature extraction and efficient computation.
-Models such as SRResNet [8] and EDSR [9] have demonstrated that deeper CNNs can progressively enhance SR performance.
-However, they suffer from the Effective Receptive Field (ERF) problem (as depicted in Fig.
-1 (a)).
-Although deeper networks theoretically have larger receptive fields, only a small portion is effectively utilized during feature extraction, which limits their ability to capture long-range dependencies.
-This becomes particularly challenging in RSIs, where diverse spatial patterns and textures overwhelm the limited modeling capacity of CNN-based methods [10, 11].
-In recent years, Transformers have gained significant traction in computer vision [14, 15], including RSISR [10, 16].
-They excel at capturing global dependencies by modeling long-range interactions through self-attention mechanisms.
-However, the quadratic complexity of the self-attention operation can become a bottleneck when processing large-scale satellites or aerial images.
-To address this, many Transformer-based models aim to reduce computational demands by limiting attention to smaller windows [14], which in turn reduces the effective receptive field (as shown in Fig.
-1 (b)) and conflicts with the goal of capturing long-range dependencies.
-Additionally, HAT [15] combines channel attention and window-based self-attention with an overlapping cross-attention module to capture long-range dependencies.
-However, although transformers can capture the global context, they struggle with computational complexity and may lose structural information during sequence transformation [17], necessitating a re-evaluation of existing methods and the development of novel approaches.
-The linear architecture of Transformers like Mamba-based models [18] and RWKV-based models [13] serves as an efficient alternative to Transformers.
-Mamba-based models leverage State Space Models (SSMs) [19] to causally capture long-range dependencies with linear computational complexity, with recent studies suggesting that they can achieve performance comparable to or even better than Transformer-based models [18, 20].
+In particular, remote sensing image super-resolution (RSISR) serves as a pivotal technique for enhancing spatial resolution to extract fine-grained details required for precise geospatial interpretation.
+However, quantitative frequency analysis highlights a fundamental imbalance in spatial frequency composition between LR and HR observations.
+As illustrated in Fig.
+1, LR images predominantly contain low-frequency components, while critical high-frequency features representing geospatial patterns are severely attenuated.
+This asymmetry is exacerbated by conventional degradation processes that disproportionately suppress high-frequency spatial information, creating reconstruction bottlenecks that standard SR methods struggle to resolve effectively.
+Traditional SR methods have been dominated by Convolutional Neural Networks (CNNs) [6, 7, 8, 9] due to their strong local feature extraction and computational efficiency.
+However, as depicted in Fig.
+2 (a), the limited effective receptive field (ERF) [10] of CNNs inherently restricts their ability to model long-range dependencies.
+This becomes particularly challenging in large-scale RSIs, where diverse spatial patterns and textures overwhelm the limited modeling capacity of CNN-based methods [11].
+While transformer-based models [12, 11, 13] alleviate the ERF limitation through global self-attention mechanisms, their quadratic computational complexity necessitates window-based self-attention that paradoxically shrinks effective receptive field coverage (as depicted in Fig.
+2 (b)), and their sequence transformation processes risk losing structural information critical for spatial consistency in RSIs [14].
+Recent advances in linear architectures address the computational constraints of Transformers through distinct approaches.
+Mamba-based models [15, 16] leverage state-space sequences (SSMs) [17] to achieve linear-complexity long-range dependency modeling, demonstrating performance comparable to Transformers.
 However, their inherent sequence modeling in SSMs still poses challenges in achieving an optimal ERF for 2D images (as depicted in Fig.
-1 (c)).
-RWKV-based models [21, 22], originally developed for natural language processing (NLP), has emerged as a potential alternative to the transformer and amba.
-RWKV improves efficiency with linear complexity for long-range dependencies and better captures local dependencies, enhancing performance on large-scale NLP and vision tasks.
-Recent studies have shown that RWKV-based models outperform Transformer-based models and linearly scaled Mamba-based models in terms of effectiveness and efficiency [22, 23].
+2 (c)), limiting spatial pattern capture in RSIs.
+Conversely, RWKV-based models combine linear attention mechanisms with localized dependency modeling, outperforming both Transformers and Mamba variants in vision tasks [18, 19, 20] and offering efficient long-range modeling with expanded ERF (as depicted in Fig.
+2 (d)).
 Despite pioneering efforts to adopt RWKV for vision tasks, the adaptation of RWKV to the field of RSISR remains unexplored.
-While CNNs, Transformers, Mamba, and RWKV networks each offer distinct advantages, the challenges of effectively managing receptive fields and addressing the ERF limitations remain open research questions in RSISR.
-Notably, a recent study [24] has shown that multi-head self-attention (MSA) and convolutional layers exhibit complementary filtering behaviors.
-Specifically, MSAs act as low-pass filters, emphasizing broader structures and smoothing noise, whereas conv layers serve as high-pass filters, focusing on fine-grained textures and edges.
-These contrasting properties make the combination of MSAs and conv layers particularly beneficial for improving model performance.
-Moreover, high-frequency details are crucial for accurately reconstructing remote sensing images, as they play a key role in preserving fine-grained information.
-However, existing approaches struggle to simultaneously model long-range dependencies while effectively capturing high-frequency details.
-Therefore, a natural question arises: Can a more efficient yet effective solution be developed to capture long-range dependencies across large-scale RSIs with convolution, which directly handles 2D spatial structures and offers computational advantages, while also efficiently reconstructing high-frequency details?
-In this paper, we propose a novel approach that integrates global representations and detailed extraction with wavelet losses to enhance the SR performance for remote sensing imagery.
+Crucially, existing approaches spanning CNNs, Transformers, Mamba, and RWKV variants universally neglect the complementary synergy between global attention mechanisms (low-pass filtering) and convolutional operations (high-pass filtering) [21], failing to synergize structural context preservation with high-frequency detail recovery essential for precise large-scale RSI reconstruction.
+This gap underscores the need for architectures that unify efficient global dependency modeling and localized texture refinement tailored to RSISR.
+In this paper, we propose a novel approach that integrates global representations and detailed extraction with wavelet losses to enhance the performance of RSISR.
 Our main contributions can be summarized as follows:
 1)
-The RWKV is introduced for remote sensing image super-resolution, along with GDSR, a novel architecture designed to effectively combine global and local information through a dual-branch structure.
+GDSR (Global-Detail Super-Resolution) is proposed to synergistically unify the RWKV and CNN within a dual-branch framework, where global contextual dependencies and localized spatial features are collaboratively refined for RSISR.
 2)
-A Global–Detail Reconstruction Module (GDRM) is proposed to efficiently integrate information from the dual branches.
+A Global–Detail Reconstruction Module (GDRM) is proposed to integrate information from the dual branches efficiently.
 By merging complementary features, this module enhances the synergy between the branches, thereby significantly improving the overall performance of RSISR tasks.
 3)
-Wavelet loss is introduced as a simple yet powerful loss function for RSISR.
-This approach effectively enhanced both the PSNR and visual quality, and demonstrated robust performance improvements across various network architectures.
-Moreover, its ease of implementation and adaptability make it a practical and highly effective solution for advancing RSISR techniques.
-We demonstrate the empirical validity of GDSR by comparing it with various competitive CNN-, GAN-, Transformer-, Mamba-, and hybrid-based models on two RSI benchmark datasets using two degradation methods showing SOTA results.
-In addition, we conducted extensive ablation experiments (as detailed in Section IV-C) to investigate the approach of global-detail integration and its effect on RSI-SR, and to examine the effect of Wavelet Loss on various model architectures.
+Dual-Group Multi-Scale Wavelet Loss is proposed to unify low-frequency structural subbands and aggregated high-frequency detail subbands through a dual-coefficient grouping strategy, enforcing cross-resolution frequency consistency via hierarchical wavelet decomposition across resolutions within a multi-scale framework, leveraging wavelet-domain frequency priors to augment spatial-domain optimization.
+We demonstrate the empirical validity of GDSR by comparing it with various competitive CNN-, GAN-, Transformer-, Mamba-, and hybrid-based models on three RSI benchmark datasets using two degradation methods showing SOTA results.
+In addition, we conducted extensive ablation experiments (as detailed in Section IV-C) to investigate the approach of global-detail integration and its effect on RSI-SR, and to examine the effect of Dual-Group Multi-Scale Wavelet Loss.
 We believe our work will advance efficient large-scale RSI-SR, improving both fidelity and processing speed, while ensuring robustness and practical applicability for real-world remote sensing tasks.
@@
 Recent advances in deep learning have greatly improved RSI-SR, with methods typically falling into CNN-based, Transformer-based, and Mamba-based categories.
-CNN-based methods, inspired by SRCNN [6], have used modules like residual connections [7], dense structures [25], and attention mechanisms [26], [27] to enhance performance.
+CNN-based methods [22, 23, 24, 25], inspired by SRCNN [6], have used modules like residual connections and attention mechanisms to enhance performance.
 However, they are limited by their inability to capture long-range dependencies, which hinders their effectiveness in large-scale RSI data.
-Transformer-based methods, leveraging self-attention (SA) [28], have shown promise in modeling global dependencies, with approaches like window-based SA [15] and recursive SA [29] attempting to address the quadratic complexity.
-However, SA remains inefficient for high-resolution images, often sacrificing global modeling capability.
-To overcome these limitations, Mamba-based methods, which offer linear complexity for global modeling, present a compelling solution.
-MambaIR [18] successfully introduced Mamba into the SR domain, while FreMamba [20] pioneers the exploration of Mamba’s potential for RSI-SR, extending it with frequency analysis.
-Due to the unidirectional nature of SSM, Mamba’s global dependency modeling still faces challenges in achieving accurate results for RSI-SR.
+Transformer-based methods [11, 26, 27] leverage self-attention (SA) to overcome CNN’s locality constraints.
+While standard SA achieves global interaction through pairwise token correlation, its quadratic complexity becomes prohibitive for high-resolution RSIs.
+Window-based SA [12] reduces computation via local window partitioning, yet critically limits cross-window dependencies required for reconstructing continuous geographic features (e.g., highway systems or river networks) spanning multiple windows.
+Mamba-based methods emerge as an alternative with linear complexity.
+MambaIR [15] successfully introduced Mamba [17] into the SR domain, while FreMamba [16] pioneers the exploration of Mamba’s potential for RSI-SR, extending it with frequency analysis.
+However, SSMs’ causal modeling framework, constrained by 1D sequence scanning of 2D spatial data, fundamentally struggles to capture omnidirectional contextual relationships required for coherent reconstruction of geometrically complex remote sensing features (e.g., roundabouts with radial roads).
+Emerging hybrid architectures combining multiple paradigms have demonstrated significant progress.
+Recent works [28, 29, 30, 31] propose innovative solutions integrating CNN-Transformer feature fusion, U-Net structures with adaptive edge recovery, and shift channel attention-enhanced Transformers, effectively addressing limitations of single-paradigm approaches while maintaining computational efficiency.
 II-B RWKV in computer vision
-Recent studies [22] indicate that RWKV can outperform both Transformers and Mamba in NLP tasks.
-Vision-RWKV [30] has been successfully applied to vision tasks, showing superior performance compared to vision Transformers with lower computational complexity.
-Several RWKV-based models, such as Diffusion-RWKV [31] for image generation and Restore-RWKV [13] for medical image restoration, have demonstrated promising results.
-However, the potential of RWKV for RSI-SR remains unexplored.
-Given that RSI exhibits diverse and complex characteristics, with texture information often subtler than in natural images, it presents unique challenges for SR tasks.
-Moreover, critical high-frequency details tend to be lost in deep learning models, emphasizing the need for a more effective approach to capture these crucial components.
-This paper pioneers the exploration of RWKV’s potential for RSI-SR by introducing a novel dual-branch architecture that combines RWKV with convolutional operations.
+Recent studies [32, 18, 19] highlight the superiority of RWKV over Transformers and Mamba in NLP tasks due to its linear computational complexity and parallelizable recurrent architecture.
+In computer vision, Vision-RWKV [33] adapts RWKV for 2D spatial processing by introducing a bidirectional WKV attention mechanism to capture global dependencies and a quad-directional token shift operation to model local contextual relationships across horizontal and vertical axes.
+These adaptations preserve the efficiency of RWKV while enabling effective feature learning in images.
+Subsequent developments further expand RWKV’s applicability.
+Diffusion-RWKV [34] integrates RWKV into diffusion models for image generation, leveraging bidirectional processing to handle patched image inputs through linear-complexity spatial aggregation, thereby eliminating the need for window-based operations.
+Restore-RWKV [20] enhances 2D dependency modeling through a recurrent WKV (Re-WKV) mechanism that integrates bidirectional and recurrent attention across multiple scan directions, alongside an omnidirectional token shift (Omni-Shift) layer to broaden context aggregation.
+However, existing implementations remain suboptimal for RSI due to distinct domain characteristics.
+Restore-RWKV’s homogeneous U-Net backbone suits anatomical medical data but falters with RSI’s extreme scale variations and heterogeneous geospatial textures.
+This paper introduces a dual-branch framework that decouples global feature extraction using RWKV from local feature extraction using CNN, directly addressing RSI-SR-specific challenges unmet by homogeneous architectures.
 II-C Wavelet Transforms in Deep Learning
-The wavelet transform (WT) [32] has been a fundamental tool for signal processing since the 1980s and has more recently been integrated into neural network architectures for various tasks.
-Several studies have explored wavelet-domain methods for image restoration and SR tasks, recognizing the need to treat low- and high-frequency features differently.
-Methods like Wavelet-SRNet [33], WDST [34] and WDRN [35] predict wavelet coefficients to enhance SR performance.
-More recently, WGSR [36] introduced wavelet subband losses in GAN-based SR models to better control artifacts.
-In addition, WTConv [37], a novel layer leveraging the wavelet transform to effectively expand the receptive field of convolutions, has been proposed as a drop-in replacement for depth-wise convolutions in CNNs, improving performance across several key computer vision tasks.
-In this paper, we propose a novel Wavelet Loss that directly optimizes the reconstruction of wavelet-domain features.
-Our approach is designed to complement SR models based on various architectures, including CNN, GAN, Transformer, and Mamba, as well as their hybrid combinations, which have demonstrated superior performance in RSISR.
-Unlike previous methods, our Wavelet Loss seamlessly integrated wavelet-based supervision into the end-to-end SR training framework to ensure effective high-frequency detail recovery.
+The wavelet transform (WT) [35] has been a fundamental signal processing tool since the 1980s and has recently been integrated into neural network architectures.
+The initial wave of deep learning integration saw wavelet-based architectures like Wavelet-SRNet [36], WDST [37] and WDRN [38] employing wavelet coefficient prediction for SR enhancement.
+Subsequent innovations deepened this synergy, with WTConv [39] introducing trainable wavelet-based layers that dynamically expand convolutional receptive fields, demonstrating cross-task adaptability in computer vision.
+The integration further evolved through adversarial frameworks, particularly WGSR [40] that embedded wavelet subband losses within GAN architectures.
+Moreover, [41] extended this strategy by integrating wavelet and RGB losses in Transformer-based models, enhancing high-frequency preservation and achieving superior visual performance on natural images compared to pixel-wise loss training.
+Notably, existing wavelet-based SR methods typically process subbands at a single scale, failing to capture the multi-scale characteristics intrinsic to RSI textures.
+This fragmented optimization approach separately constrains low-frequency structural components and high-frequency directional details without explicit cross-scale coordination, leading to inconsistent reconstruction of hierarchical spatial patterns.
+In this paper, we propose a novel Dual-Group Multi-Scale Wavelet Loss to enforce cross-resolution frequency consistency through explicit approximation subband grouping and multi-scale high-frequency aggregation (horizontal/vertical/diagonal).
 III Methodology
-In this section, we first introduce the overall structure of the proposed GDSR and then described three important modules in GDSR, namely RGEG, RDEG, and GDRM.
+In this section, we first introduce the overall structure of the proposed GDSR and then describe three important modules in GDSR, namely RGEG, RDEG, and GDRM.
 Finally, we detail our designed loss function.
@@
 As illustrated in Fig.
-2, GDSR begins by extracting the shallow feature from the LR input using a convolutional layer.
+3, GDSR begins by extracting the shallow feature from the LR input using a convolutional layer.
 It then derives the deep features and through parallel branches: global extraction branch and details extraction branch, respectively.
 These features are subsequently fused using the GDRM.
-A global residual connection is then applied to combine the low-level and deep features, and finally, the pixel-shuffle method is employed for upsampling to generate the SR output.
+A global residual connection is then applied to combine the low-level and deep features, and finally, the pixel-shuffle method [42] is employed for upsampling to generate the SR output.
 The SR process of GDSR can be mathematically described as:
@@
 As illustrated in Fig.
-2, the proposed RGEG consists of multiple Residual RWKV Blocks (R-RWKVBs).
+3, the proposed RGEG consists of multiple Residual RWKV Blocks (R-RWKVBs).
 The spatial mix (sm) module is designed to establish long-range dependencies among tokens across the spatial dimension.
@@
 where , , and are linear projection matrices, and is the squared ReLU activation function, which enhances nonlinearity.
-Notably, the transformation from to to involves a multi-layer perception (MLP) consisting of , , and , facilitating channel-wise feature fusion.
+Notably, the transformation from to to involves a multi-layer perception consisting of , , and , facilitating channel-wise feature fusion.
 The final channel mix output is obtained through:
@@
 As illustrated in Fig.
-3, the module begins by applying a permute operation to the input feature map, reorganizing it into three distinct formats, each capturing specific spatial relationships: Height-Width (HW), Channel-Width (CW), and Height-Channel (HC).
+4, the module begins by applying a permute operation to the input feature map, reorganizing it into three distinct formats, each capturing specific spatial relationships: Height-Width (HW), Channel-Width (CW), and Height-Channel (HC).
 Spatial Attention [43] is subsequently applied to each permuted representation independently, enabling the modeling of dependencies along these dimensions.
@@
 As illustrated in Fig.
-4, the GDRM receives two input feature maps: one capturing local details from the RDEG branch and another representing global contextual information from the RGEG branch.
+5, the GDRM receives two input feature maps: one capturing local details from the RDEG branch and another representing global contextual information from the RGEG branch.
 The GDRM then processes these feature maps through a series of operations, which can be expressed as:
@@
 […]
-III-E Wavelet Loss
-Due to the unique characteristics of remote sensing imagery, preserving fine-grained details across multiple scales is critical for enhancing spatial resolution while preserving fine-grained spatial details and structural consistency.
-To address this challenge, we introduce an integrable wavelet-based loss function designed for versatile and effective application across diverse SR models, leveraging the Haar wavelet transform to capture high-frequency details essential for visually and quantitatively improved RSI-SR results.
+III-E Dual-Group Multi-Scale Wavelet Loss
+The proposed wavelet loss construction follows the paradigm established in prior studies [40, 41].
 As illustrated in Fig.
-5, the SWT generates one low-frequency (LF) subband, LL, representing the coarse structure of the image, and three high-frequency (HF) subbands, LH, HL, and HH, which encode horizontal, vertical, and diagonal details, respectively.
-As illustrated in Fig.
-6, the wavelet loss function is formulated as the distance between the corresponding Haar wavelet subbands of the generated SR image and the HR image.
-The total wavelet loss is expressed as:
-where and are the HR and SR images, denotes the Haar wavelet transform, indexes the subbands, and are scaling factors that balance the importance of each subband.
-The total wavelet loss is averaged over a minibatch size represented by .
-By emphasizing high-frequency subbands, the loss effectively enhances edge and texture preservation, while the low-frequency subband ensures structural consistency.
+6, initially, both the HR image and SR image are transformed into the YCbCr color space.
+The stationary wavelet transform (SWT) is subsequently applied exclusively to the luminance (Y) channel, as human visual perception demonstrates higher sensitivity to structural details in this component.
+Through SWT decomposition, one low-frequency (LF) subband (LL) and three high-frequency (HF) subbands (LH, HL, HH) are generated.
+The LL subband encapsulates coarse structural information, while the HF components encode directional details: LH captures horizontal features, HL preserves vertical characteristics, and HH retains diagonal patterns.
+These HF subbands are concatenated along the channel dimension to form a composite HF representation , whereas the LL subband is retained as .
+The wavelet loss function is formulated as the distance between the corresponding subbands of the generated SR image and the HR image. The total wavelet loss is expressed as:
+where denotes three spatial resolutions, , and , represent the LF/HF subbands of the SR/HR images at scale , are resolution-specific weights satisfying (Specific settings for can be found in Table VI, but generally, arbitrary can be used), and , control the global contributions of LF/HF components.
 The overall loss for the training with wavelet loss is given by
 IV Experiment
+IV-A Datasets and Evaluation
+This paper reports the performance of SR on three RSI datasets, including the publicly available AID [44], UCMerced [45], and the custom RSSRD-QH dataset.
+The AID dataset consists of 30 categories, with the data in each category randomly divided into two equal parts for training and testing.
+Additionally, 20% of the training data was reserved as a validation set, resulting in 4,000 training images, 1,000 validation images, and 5,000 test images.
+Similarly, the UCMerced dataset, comprising 21 categories with a total of 2,100 images, was processed identically, yielding 945 training images, 105 validation images (10% of the training data), and 1,050 test images.
+The RSSRD-QH dataset is located in representative watershed units within the Sanjiangyuan region and the Qinghai Lake surrounding area.
+The geographical distribution of the counties covered by the dataset, including their longitude and latitude ranges, is summarized in Table I.
+The dataset was collected using unmanned aerial vehicles, with flight altitudes carefully controlled between 30 and 50 meters.
+The average spatial resolution of the captured imagery is 0.01 meters, providing comprehensive coverage of various landscape elements across different elevation gradients, including terraces, floodplains, gentle slopes, and steep slopes within these watershed units.
+From each Watershed, we randomly selected 50% of the images for the training set and 50% for the test set, with 20% of the training set randomly selected for validation.
+This resulted in a training set of 2,242 images, a validation set of 561 images, and a test set of 2,804 images.
+The HR images in the AID and RSSRD-QH datasets were 600 × 600 pixels, while those in the UCMerced dataset were 256 × 256 pixels.
+The proposed method, along with other competing methods, was evaluated on the test datasets using two classic full-reference metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [46].
+Additionally, the LPIPS [47] metric was used to capture the perceptual quality of the images.
+Notably, PSNR and SSIM were calculated based on the luminance channel (Y) of the YCbCr color space, while LPIPS evaluated the perceptual similarity between images by extracting features from a pre-trained deep network, specifically the AlexNet [48], which was employed in this study.
+IV-B Degradation Model and Implementation Details
+We employed both bicubic downsampling and a comprehensive degradation model (CDM) to simulate LR images during training.
+Based on existing blind SR methods [49, 50] and the characteristics of remote sensing sensors [51, 52], the CDM for LR image synthesis included isotropic Gaussian blur, motion blur, scaling with different interpolation methods, additive Gaussian noise, and JPEG compression noise, all applied in random order.
+Our GDSR performed deep feature exploration through 4 RGEGs, RDEGs, and GDRMs.
+Each RGEG consisted of 6 R-RWKVBs, and each RDEG contains 12 RCBs.
+Empirically, we set the internal channel dimension to = 96.
+For the GDSR_TC and GDSR_MC, the RGEG branches correspond to the RTSB and ResidualGroup of SwinIR [12] and MambaIR [15], respectively, with all other settings remaining consistent with GDSR.
+During training, 64 × 64 patches were randomly cropped from LR RSIs, along with their HR counterparts scaled by the target scale factor (×2, ×3, ×4).
+The batch size was set to 16.
+The initial learning rate was set to and optimized using the Adam optimizer, where = 0.9 and = 0.99, with a step learning rate scheduler halving the rate at the 100th epoch over a total of 200 training epochs.
+All SR models were implemented in the PyTorch framework.
+The FLOPs results were calculated using an input tensor of size 1 × 3 × 160 × 160, and the computations were performed using torch-operation-counter 111https://github.com/SamirMoustafa/torch-operation-counter.
+The inference times were tested on 300 random images with a size of 3 × 160 × 160.
+All experiments were conducted on an NVIDIA A800 80GB GPU.
 | County | Longitude (°E) | Latitude (°N) |
@@
 | Tianjun | 98.409 to 98.624 | 37.45 to 37.4855 |
-IV-A Datasets and Evaluation
-This paper reports the performance of SR on two remote sensing image datasets, including the publicly available AID dataset [39] and the custom RSSRD-QH dataset.
-The AID dataset consists of 30 categories.
-For each category, the data was randomly split into two equal parts, one for training and the other for testing.
-Additionally, 20% of the training data was further separated and used as a validation set.
-Specifically, the training set contained 4,000 images, the validation set 1,000 images, and the test set 5,000 images.
-The RSSRD-QH dataset is located in representative watershed units within the Sanjiangyuan region and the Qinghai Lake surrounding area.
-The geographical distribution of the counties covered by the dataset, including their longitude and latitude ranges, is summarized in Table I.
-The dataset was collected using unmanned aerial vehicles, with flight altitudes carefully controlled between 30 and 50 meters.
-The average spatial resolution of the captured imagery is 0.01 meters, providing comprehensive coverage of various landscape elements across different elevation gradients, including terraces, floodplains, gentle slopes, and steep slopes within these watershed units.
-From each Watershed, we randomly selected 50% of the images for the training set and 50% for the test set, with 20% of the training set randomly selected for validation.
-This resulted in a training set of 2,242 images, a validation set of 561 images, and a test set of 2,804 images.
-The HR images in both datasets were 600 × 600 pixels, and the corresponding LR images were 200 × 200 pixels.
-The proposed method, along with other competing methods, was evaluated on the test datasets using two classic full-reference metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [40].
-Additionally, the LPIPS [41] metric was used to capture the perceptual quality of the images.
-Notably, PSNR and SSIM were calculated based on the luminance channel (Y) of the YCbCr color space, while LPIPS evaluated the perceptual similarity between images by extracting features from a pre-trained deep network, specifically the AlexNet model [42], which was employed in this study.
-IV-B Degradation Model and Implementation Details
-We employed both bicubic downsampling and a comprehensive degradation model (CDM) to simulate LR images during training.
-Based on existing blind SR methods [43, 44] and the characteristics of remote sensing sensors [45], the CDM for LR image synthesis included isotropic Gaussian blur, motion blur, scaling with different interpolation methods, additive Gaussian noise, and JPEG compression noise, all applied in random order.
-Our GDSR performed deep feature exploration through 4 RGEGs, RDEGs, and GDRMs.
-Each RGEG consisted of 6 R-RWKVBs and each RDEG contains 12 RCBs.
-Empirically, we set the internal channel dimension to = 96.
-For the GDSR_TC and GDSR_MC, the RGEG branches correspond to the RTSB and ResidualGroup of SwinIR [14] and MambaIR [18], respectively, with all other settings remaining consistent with GDSR.
-During training, each HR image was randomly cropped to a size of 192 × 192 pixels, with the corresponding LR image size being 64 × 64 pixels.
-The batch size was set to 16, and the Adam optimizer was used with = 0.9 and = 0.999.
-The initial learning rate was set to , which was halved at the 100th epoch, with a total of 200 training epochs.
-All SR models were implemented in the PyTorch framework.
-The FLOPs results were calculated using an input tensor of size 1 × 3 × 240 × 240, and the computations were performed using torch-operation-counter [46].
-The inference times were tested on 300 random images with a size of 3 × 240 × 240.
-All experiments were conducted on an NVIDIA A800 80GB GPU.
-RCBs | PSNR/SSIM | LPIPS | #Param. | FLOPs | FPS |
-1 | 28.96/0.7658 | 0.3514 | 5.86M | 340.25G | 7.3 |
-4 | 28.96/0.7657 | 0.3518 | 7.85M | 455.31G | 7.0 |
-8 | 29.00/0.7677 | 0.3499 | 10.51M | 608.73G | 6.7 |
-12 | 29.03/0.7686 | 0.3476 | 13.17M | 762.14G | 6.4 |
-16 | 29.01/0.7680 | 0.3480 | 15.83M | 915.55G | 6.1 |
-20 | 29.03/0.7688 | 0.3477 | 18.48M | 1068.97G | 5.8 |
 IV-C Ablation Study
 In this section, we discuss the proposed GDSR in depth by investigating the effect of its major components and their variants.
+All experiments were conducted under the ×3 super-resolution configuration with the loss function in the AID [44] dataset.
 IV-C1 Effect of Key Modules
-(a) Effect of RDEG. To investigate the effect of RDEG on the model performance, we conducted ablation experiments by varying the number of RCBs in RDEG.
-The experimental results, presented in Tables II and III, indicated that the model achieved a PSNR of 28.94 dB and an SSIM of 0.7652 when only the RGEG branch was utilized without incorporating RDEG.
-Adding a single RCB significantly improved the model’s performance, increasing the PSNR to 28.96 dB and the SSIM to 0.7658.
-This improvement highlighted the RCB’s capability to capture critical information, thereby enhancing reconstruction quality.
-As the number of RCBs increased, further improvements in PSNR and SSIM were observed.
-For instance, increasing the RCB count from 1 to 20 raised the PSNR from 28.96 dB to 29.03 dB and the SSIM from 0.7658 to 0.7688.
-However, the rate of improvement diminished as the RCB count increased, and beyond a certain threshold, performance even began to decline.
-This trend suggested that while adding more RCBs initially benefited the model’s output, the marginal gains decreased and eventually plateaued.
-Additionally, increasing the number of RCBs led to a linear rise in computational complexity, as evidenced by the increase in parameters, FLOPs, and a decrease in FPS.
-This highlighted the trade-off between performance gains and computational efficiency.
-Therefore, determining the optimal number of RCBs is crucial in practical applications, especially in scenarios where computational efficiency is a critical factor.
+(a) Effect of RDEG and RGEG.
+Ablation studies were conducted to evaluate the impact of varying RCBs in RDEG and R-RWKVBs in RGEG on model performance.
+As shown in Tables II and III, the baseline model using only the RGEG branch yielded a PSNR of 28.94 dB and SSIM of 0.7652.
+Introducing a single RCB to RDEG improved performance to 28.96 dB PSNR and 0.7658 SSIM, validating the RCB’s ability to enhance reconstruction quality.
+Scaling the RCB count to 12 further increased PSNR to 29.03 dB and SSIM to 0.7686, with a computational footprint of 13.17M parameters and 338.73G FLOPs.
+However, beyond 12 RCBs, performance plateaued (e.g., 29.01 dB with 16 RCBs) or exhibited instability (e.g., 29.03 dB with 20 RCBs), while computational costs grew linearly, reducing FPS from 14.8 to 12.6.
+This suggests diminishing returns from excessive RCB stacking, necessitating a balance between accuracy and efficiency.
+Similarly, experiments on RGEG’s R-RWKVBs revealed analogous trends.
+Fixing RCBs at 12 and increasing R-RWKVBs from 1 to 8 marginally improved PSNR and SSIM but incurred significant computational overhead, with parameters rising from 10.62M to 14.19M and FPS dropping sharply from 40.9 to 10.5.
+Notably, the joint configuration of 12 RCBs and 6 R-RWKVBs achieved optimal trade-offs, outperforming isolated branch evaluations (as shown in Table III).
+These results underscore the complementary roles of RDEG and RGEG in feature refinement while emphasizing the necessity of constrained component counts to mitigate redundancy and computational bloat in practical deployments.
+| RCBs | R-RWKVBs | PSNR | SSIM | #Param. | FLOPs | FPS | |
+| 1 |
+28.96 | 0.7658 | 5.86M | 151.23G | 15.7 | |
+| 4 | 28.96 | 0.7657 | 7.85M | 202.37G | 15.2 | ||
+| 8 | 29.00 | 0.7677 | 10.51M | 270.55G | 14.5 | ||
+| 12 | 29.03 | 0.7686 | 13.17M | 338.73G | 14.8 | ||
+| 16 | 29.01 | 0.7680 | 15.83M | 406.92G | 13.3 | ||
+| 20 | 29.03 | 0.7688 | 18.48M | 475.10G | 12.6 | ||
+1 | 29.00 | 0.7676 | 10.62M | 272.97G | 40.9 | |
+| 2 | 29.01 | 0.7678 | 11.13M | 286.12G | 29.4 | ||
+| 4 | 29.01 | 0.7679 | 12.15M | 312.43G | 16.2 | ||
+| 8 | 29.01 | 0.7679 | 14.19M | 365.04G | 10.5 |
 (b) Effect of GDRM.
-To evaluate the effectiveness of the GDRM, we conducted ablation experiments separately using the RDEG and RGEG branches.
-The experimental results were presented in Table III.
-The comparisons demonstrated that introducing the GDRM significantly improved performance compared to single-branch models.
-The GDSR model achieved superior PSNR, SSIM, and LPIPS values, indicating that GDSR effectively integrated the local detail information from RDEG and the global contextual information from RGEG.
-Additionally, the incorporation of PSAM enabled GDRM to fuse the information from both branches more effectively.
-(c) Visualization of GDRM.
-Fig.
-7 visualized the features of the RDEG, RGEG branches, and GDRM.
-We observed that the feature maps from RDEG and RGEG were complementary, while GDRM combined features from both branches.
-This demonstrated that, within the dual-branch structure of GDSR, each branch focused on different features, and fusing these features enabled effective image reconstruction.
+Ablation experiments in Table III demonstrated that the dual-branch GDSR model outperformed single-branch alternatives, achieving superior PSNR, SSIM, and LPIPS values.
+This performance gain was attributed to the GDRM, which systematically integrated dual-branch features through its PSAM and adaptive fusion mechanism (Eqs.
+14 and 15).
+From a spatial-frequency duality perspective, RGEG’s global attention mechanisms implicitly exhibited low-pass filtering characteristics to maintain structural continuity, while RDEG’s convolutional operations inherently emphasized high-frequency components for local detail enhancement [21].
+As visualized in Fig.
+7, the dual-branch structure of GDSR demonstrated complementary feature specialization with GDRM dynamically fusing both to enable efficient reconstruction.
+Visualizations in Fig.
+8 revealed that RGEG activated distributed pixels across broader regions, whereas RDEG focused on local patterns.
+The F2D strategy further enhanced this synergy by routing refined features exclusively to RDEG, prioritizing localized refinement over redundant global propagation.
+This design addressed a critical theoretical limitation: while RGEG aggregated broader pixel information, effective reconstruction required targeted utilization of spatial relationships within receptive fields.
+F2D’s spectral separation strategy specifically resolved conflicts observed in F2B, where dual-path feedback introduced frequency interference, as global attention diluted high-frequency details while convolutions struggled to recover them.
+The quantitative results confirmed that GDSR’s dual-branch framework, coupled with GDRM’s reconstruction, systematically unified heterogeneous features to achieve both perceptual and pixel-wise accuracy.
+This validated the theoretical rationale that explicitly decoupling global (low-frequency) and local (high-frequency) processing paths, followed by attention-driven fusion, was essential for overcoming the limitations of conventional single-path architectures.
 | Method | PSNR (dB) | SSIM | LPIPS |
@@
 | F2D(GDSR) | 29.03 | 0.7686 | 0.3476 |
-Depth | PSNR/SSIM | LPIPS | #Param. | FLOPs | FPS |
-1 | 28.89/0.7633 | 0.3565 | 3.65M | 211.64G | 24.6 |
-2 | 28.97/0.7663 | 0.3514 | 6.82M | 395.14G | 12.6 |
-4 | 29.03/0.7686 | 0.3476 | 13.17M | 762.14G | 6.4 |
-6 | 29.03/0.7686 | 0.3463 | 19.52M | 1129.14G | 4.3 |
-8 | 29.04/0.7692 | 0.3456 | 25.86M | 1496.14G | 3.2 |
-10 | 29.05/0.7695 | 0.3450 | 32.21M | 1863.13G | 2.5 |
-Methods | Venue | AID [39] | AID_CDM [39] | RSSRD-QH | RSSRD-QH_CDM | ||||||||
-PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | ||
-Bicubic | - | 27.89 | 0.7238 | 0.4400 | 23.15 | 0.5238 | 0.6911 | 26.45 | 0.6447 | 0.4243 | 22.34 | 0.3525 | 0.6993 |
-SRCNN [6] | CVPR’2014 | 28.22 | 0.7363 | 0.3973 | 23.38 | 0.5455 | 0.6569 | 26.54 | 0.6512 | 0.4241 | 22.44 | 0.3588 | 0.7523 |
-SRGAN [8] | CVPR’2017 | 28.80 | 0.7597 | 0.3601 | 23.78 | 0.5648 | 0.6314 | 26.68 | 0.6586 | 0.4170 | 22.64 | 0.3759 | 0.7421 |
-EDSR [9] | CVPR’2017 | 28.94 | 0.7653 | 0.3513 | 23.91 | 0.5698 | 0.6139 | 26.70 | 0.6598 | 0.4141 | 22.74 | 0.3817 | 0.7283 |
-SwinIR [14] | CVPR’2021 | 28.95 | 0.7658 | 0.3512 | 23.89 | 0.5705 | 0.6169 | 26.70 | 0.6597 | 0.4133 | 22.72 | 0.3780 | 0.7432 |
-HAT [15] | CVPR’2023 | 28.99 | 0.7671 | 0.3479 | 23.99 | 0.5737 | 0.6080 | 26.71 | 0.6602 | 0.4111 | 22.79 | 0.3823 | 0.7312 |
-MambaIR [18] | ECCV’2024 | 29.02 | 0.7684 | 0.3487 | 24.07 | 0.5763 | 0.6000 | 26.72 | 0.6607 | 0.4124 | 22.84 | 0.3835 | 0.7387 |
-ConvFormerSR [16] | TGRS’2024 | 28.90 | 0.7637 | 0.3516 | 23.97 | 0.5710 | 0.6118 | 26.69 | 0.6603 | 0.4090 | 22.77 | 0.3779 | 0.7365 |
-FreMamba [20] | TMM’2024 | 28.96 | 0.7672 | 0.3410 | 23.89 | 0.5692 | 0.6148 | 26.70 | 0.6601 | 0.4112 | 22.60 | 0.3660 | 0.7242 |
-GDSR_TC (Ours) | - | 28.96 | 0.7660 | 0.3500 | 24.01 | 0.5717 | 0.6126 | 26.71 | 0.6604 | 0.4144 | 22.79 | 0.3776 | 0.7401 |
-GDSR_MC (Ours) | - | 28.97 | 0.7668 | 0.3472 | 24.03 | 0.5734 | 0.6080 | 26.71 | 0.6602 | 0.4107 | 22.84 | 0.3819 | 0.7327 |
-GDSR (Ours) | - | 29.03 | 0.7686 | 0.3476 | 24.10 | 0.5770 | 0.5986 | 26.72 | 0.6612 | 0.4106 | 22.86 | 0.3861 | 0.7268 |
-GDSR_L (Ours) | - | 29.04 | 0.7692 | 0.3456 | 24.15 | 0.5795 | 0.5936 | 26.73 | 0.6615 | 0.4098 | 22.89 | 0.3875 | 0.7253 |
-Methods | Venue | AID [39] | AID_CDM [39] | RSSRD-QH | RSSRD-QH_CDM | ||||||||
-PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | ||
-SRCNN [6] | CVPR’2014 | 28.25 | 0.7369 | 0.3998 | 23.38 | 0.5456 | 0.6609 | 26.55 | 0.6498 | 0.4221 | 22.44 | 0.3590 | 0.7485 |
-SRGAN [8] | CVPR’2017 | 28.82 | 0.7605 | 0.3672 | 23.80 | 0.5652 | 0.6458 | 26.69 | 0.6581 | 0.4183 | 22.65 | 0.3760 | 0.7456 |
-EDSR [9] | CVPR’2017 | 28.96 | 0.7658 | 0.3524 | 23.94 | 0.5711 | 0.6233 | 26.70 | 0.6596 | 0.4128 | 22.76 | 0.3821 | 0.7357 |
-SwinIR [14] | CVPR’2021 | 28.98 | 0.7668 | 0.3533 | 23.90 | 0.5704 | 0.6301 | 26.71 | 0.6592 | 0.4147 | 22.73 | 0.3781 | 0.7488 |
-HAT [15] | CVPR’2023 | 29.04 | 0.7692 | 0.3483 | 24.01 | 0.5736 | 0.6183 | 26.72 | 0.6597 | 0.4137 | 22.81 | 0.3829 | 0.7348 |
-MambaIR [18] | ECCV’2024 | 29.04 | 0.7690 | 0.3521 | 24.08 | 0.5767 | 0.6114 | 26.72 | 0.6598 | 0.4134 | 22.85 | 0.3842 | 0.7388 |
-ConvFormerSR [16] | TGRS’2024 | 28.93 | 0.7647 | 0.3524 | 23.99 | 0.5707 | 0.6281 | 26.71 | 0.6591 | 0.4141 | 22.77 | 0.3770 | 0.7427 |
-FreMamba [20] | TMM’2024 | 28.94 | 0.7673 | 0.3397 | 23.89 | 0.5682 | 0.6318 | 26.70 | 0.6599 | 0.4130 | 22.64 | 0.3698 | 0.7130 |
-GDSR_TC (Ours) | - | 28.99 | 0.7670 | 0.3522 | 24.02 | 0.5719 | 0.6237 | 26.72 | 0.6595 | 0.4162 | 22.79 | 0.3776 | 0.7437 |
-GDSR_MC (Ours) | - | 29.00 | 0.7675 | 0.3512 | 24.06 | 0.5742 | 0.6159 | 26.72 | 0.6591 | 0.4122 | 22.84 | 0.3815 | 0.7395 |
-GDSR (Ours) | - | 29.04 | 0.7689 | 0.3522 | 24.12 | 0.5774 | 0.6121 | 26.73 | 0.6603 | 0.4116 | 22.88 | 0.3862 | 0.7313 |
-GDSR_L (Ours) | - | 29.06 | 0.7694 | 0.3499 | 24.16 | 0.5791 | 0.6069 | 26.73 | 0.6603 | 0.4119 | 22.90 | 0.3875 | 0.7314 |
-Fig.
-8 provided visual comparisons under different branches and reconstruction strategies.
-It was evident that the RGEG branch had a significantly larger receptive field than the RDEG branch and activated more widely distributed pixels for image reconstruction.
-As a result, the RGEG branch achieved better PSNR and SSIM values.
-Two Global-Detail reconstruction strategies, F2B and F2D, yielded different results.
-F2B, which sent the reconstructed features back to both RGEG and RDEG branches, largely retained the wide receptive field benefits of the RGEG branch and similar DI values.
-However, its SR results were comparable to those of the single RGEG branch.
-In contrast, F2D, which sent the reconstructed features only back to the RDEG branch, achieved DI values between those of the RGEG and RDEG branches.
-The SR results of the F2D strategy showed that it correctly reconstructed textures and delivered more desirable outcomes.
-This finding aligned with previous studies, which suggested that while expanding the model’s receptive field allowed consideration of a broader range of pixels, it could also lead to incorrect texture reconstruction.
-This indicated that simply leveraging information from more pixels was insufficient for accurate reconstruction; effectively utilizing pixel information within the receptive field was crucial.
-As shown in the LAM Attribution map in Fig.
-8, the F2D strategy, compared to F2B and the RGEG branch, demonstrated a greater contribution of pixels within the corresponding regions to the SR results.
-In contrast, the distant pixels captured by the wide receptive field of the RGEG branch (e.g., pixels in the upper part of the figure) contributed less to the SR results.
-This suggested that F2D, by routing features exclusively back to the RDEG branch, enabled the model to utilize information more efficiently and reconstruct accurate textures by correctly leveraging the pixels within the receptive field.
+(c) Effect of RWKV.
+To investigate the effect of RWKV, ablation experiments were conducted by replacing RWKV with Transformer [12] (GDSR_TC) and Mamba [15] (GDSR_MC) while keeping other components unchanged.
+As shown in Table IV, the RWKV-based GDSR achieves the best performance in terms of PSNR, SSIM, and LPIPS, outperforming both Transformer and Mamba variants.
+Notably, RWKV demonstrates superior computational efficiency with 14.8 FPS, significantly exceeding GDSR_TC and GDSR_MC.
+While maintaining moderate parameter size and FLOPs, RWKV achieves an optimal balance between reconstruction quality and computational cost.
+This validates RWKV’s effectiveness in capturing long-range dependencies while mitigating the quadratic complexity limitations of conventional attention mechanisms.
 (d) Effect of Network Depth on Performance.
 To investigate the effect of network depth on the performance of GDSR, we varied the shared depth of both the RDEG and RGEG branches while keeping other design parameters constant.
-As illustrated in Table IV, the experimental outcomes consistently demonstrated an enhancement in performance with deeper networks.
+As illustrated in Table V, the experimental outcomes consistently demonstrated an enhancement in performance with deeper networks.
 Moreover, an evaluation of the trade-offs in terms of computational complexity was conducted.
@@
 These results supported the effectiveness of GDSR’s dual-branch design, where the shared depth allowed for both high-quality super-resolution and manageable computational overhead.
+| Method | PSNR/SSIM | LPIPS | #Param. | FLOPs | FPS |
+| GDSR_TC | 28.89/0.7633 | 0.3565 | 12.83M | 337.26G | 7.6 |
+| GDSR_MC | 28.97/0.7663 | 0.3514 | 13.99M | 350.35G | 9.1 |
+| GDSR | 29.03/0.7686 | 0.3476 | 13.17M | 338.73G | 14.8 |
+| Depth | PSNR/SSIM | LPIPS | #Param. | FLOPs | FPS |
+| 1 | 28.89/0.7633 | 0.3565 | 3.65M | 94.06G | 53.0 |
+| 2 | 28.97/0.7663 | 0.3514 | 6.82M | 175.62G | 27.3 |
+| 4 | 29.03/0.7686 | 0.3476 | 13.17M | 338.73G | 14.8 |
+| 6 | 29.03/0.7686 | 0.3463 | 19.52M | 501.85G | 9.3 |
+| 8 | 29.04/0.7692 | 0.3456 | 25.86M | 664.96G | 7.0 |
+| 10 | 29.05/0.7695 | 0.3450 | 32.21M | 828.07G | 5.6 |
+IV-C2 Effect of Wavelet Loss
+Ablation experiments were conducted across multiple configurations to evaluate the effectiveness of incorporating wavelet loss in training.
+Quantitative results (Table VI) demonstrate that our full configuration achieves the highest PSNR and SSIM while attaining the lowest LPIPS, outperforming both baselines and the dual-group-only variant.
+Visual evidence in Fig.
+9 further corroborates these findings, the inclusion of the proposed Dual-Group Multi-Scale Wavelet Loss significantly improved the reconstruction, particularly in the high-frequency components.
+This improvement is attributed to the synergistic integration of dual-coefficient grouping strategy and cross-resolution frequency consistency, which collectively overcome the rigidity of fixed-scale wavelet methods that process subbands in isolation.
+These results highlight Wavelet Loss’s role in guiding the model to recover intricate structures and accurately reconstruct complex textures, thereby enhancing overall image quality.
+However, as evidenced in Table VI, all wavelet loss variants exhibit higher LPIPS values compared to the baseline, indicating an inherent challenge in reconciling frequency-domain fidelity with perceptual authenticity.
+Specifically, the uniform coefficient configuration [41] yields the worst LPIPS due to its undifferentiated weighting.
+Our dual-group multi-scale design strategically mitigates this tradeoff.
+While the proposed method’s LPIPS remains marginally higher than , it achieves the lowest LPIPS among wavelet-based approaches.
+The resultant configuration delivers superior PSNR/SSIM with minimal perceptual compromise, demonstrating that while frequency-domain constraints inherently elevate LPIPS, our method optimally balances this tradeoff with substantial PSNR/SSIM gains.
+| Method | PSNR (dB) | SSIM | LPIPS |
+| loss | 29.0329 | 0.7686 | 0.3476 |
+| Baseline [41]: Uniform coefficients | 29.0182 | 0.7676 | 0.3565 |
+| Baseline [41]: Asymmetric coefficients | 29.0413 | 0.7687 | 0.3534 |
+| Proposed: Dual-Group | 29.0468 | 0.7690 | 0.3535 |
+| Proposed: Dual-Group Multi-Scale | 29.0514 | 0.7694 | 0.3519 |
+| Methods | Type | AID x2 | AID x3 | AID x4 | |||||||||
+| PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |||||
+| Bicubic |
+30.48 | 0.8269 | 0.2769 | 27.89 | 0.7238 | 0.4400 | 26.43 | 0.6541 | 0.5640 | |||
+| SRCNN [6] | 30.87 | 0.8352 | 0.2711 | 28.22 | 0.7363 | 0.3973 | 26.68 | 0.6679 | 0.4862 | ||||
+| EDSR [9] | 31.70 | 0.8560 | 0.2416 | 28.94 | 0.7653 | 0.3513 | 27.30 | 0.6994 | 0.4381 | ||||
+| SRGAN [8] | 31.57 | 0.8523 | 0.2481 | 28.80 | 0.7597 | 0.3601 | 27.20 | 0.6939 | 0.4466 | ||||
+| SwinIR [12] | 31.71 | 0.8567 | 0.2370 | 28.95 | 0.7658 | 0.3512 | 27.32 | 0.6993 | 0.4373 | ||||
+| HAT [13] | 31.77 | 0.8579 | 0.2351 | 28.99 | 0.7671 | 0.3479 | 27.34 | 0.7004 | 0.4342 | ||||
+| HAUNet [25] | 31.48 | 0.8507 | 0.2503 | 28.75 | 0.7580 | 0.3657 | 27.15 | 0.6913 | 0.4523 | ||||
+| SPT [27] | 31.61 | 0.8547 | 0.2391 | 28.86 | 0.7620 | 0.3553 | 27.26 | 0.6962 | 0.4418 | ||||
+| ConvFormerSR [28] | 31.64 | 0.8551 | 0.2298 | 28.90 | 0.7637 | 0.3516 | 27.26 | 0.6964 | 0.4396 | ||||
+| MambaIR [15] | 31.78 | 0.8583 | 0.2354 | 29.02 | 0.7684 | 0.3487 | 27.37 | 0.7019 | 0.4348 | ||||
+| FreMamba [16] | 31.62 | 0.8558 | 0.2215 | 28.96 | 0.7672 | 0.3410 | 27.32 | 0.7010 | 0.4255 | ||||
+| GDSR (Ours) | 31.81 | 0.8592 | 0.2366 | 29.05 | 0.7694 | 0.3519 | 27.39 | 0.7022 | 0.4376 | ||||
+| Bicubic |
+24.11 | 0.5545 | 0.6046 | 23.15 | 0.5238 | 0.6911 | 22.54 | 0.5091 | 0.7543 | |||
+| SRCNN [6] | 24.44 | 0.5854 | 0.5836 | 23.38 | 0.5455 | 0.6569 | 22.70 | 0.5245 | 0.7100 | ||||
+| EDSR [9] | 24.90 | 0.6153 | 0.5341 | 23.91 | 0.5698 | 0.6139 | 23.16 | 0.5444 | 0.6753 | ||||
+| SRGAN [8] | 24.78 | 0.6070 | 0.5609 | 23.78 | 0.5648 | 0.6314 | 23.02 | 0.5396 | 0.6909 | ||||
+| SwinIR [12] | 24.95 | 0.6144 | 0.5353 | 23.89 | 0.5705 | 0.6169 | 23.17 | 0.5452 | 0.6737 | ||||
+| HAT [13] | 25.06 | 0.6183 | 0.5276 | 23.99 | 0.5737 | 0.6080 | 23.26 | 0.5475 | 0.6671 | ||||
+| HAUNet [25] | 25.03 | 0.6169 | 0.5287 | 23.90 | 0.5692 | 0.6185 | 23.19 | 0.5441 | 0.6783 | ||||
+| SPT [27] | 24.86 | 0.6086 | 0.5457 | 23.88 | 0.5679 | 0.6257 | 23.20 | 0.5447 | 0.6749 | ||||
+| ConvFormerSR [28] | 25.04 | 0.6137 | 0.5378 | 23.97 | 0.5710 | 0.6118 | 23.23 | 0.5448 | 0.6742 | ||||
+| MambaIR [15] | 25.17 | 0.6230 | 0.5120 | 24.07 | 0.5763 | 0.6000 | 23.32 | 0.5497 | 0.6590 | ||||
+| FreMamba [16] | 25.00 | 0.6227 | 0.4977 | 23.89 | 0.5692 | 0.6148 | 23.06 | 0.5403 | 0.6735 | ||||
+| GDSR (Ours) | 25.28 | 0.6256 | 0.5225 | 24.12 | 0.5771 | 0.6106 | 23.39 | 0.5510 | 0.6664 | ||||
+| UCMerced x2 | UCMerced x3 | UCMerced x4 | |||||||||||
+| Bicubic |
+31.10 | 0.8892 | 0.2013 | 27.59 | 0.7731 | 0.3548 | 25.73 | 0.6813 | 0.4838 | |||
+| SRCNN [6] | 32.33 | 0.9102 | 0.1635 | 28.48 | 0.8028 | 0.2919 | 26.39 | 0.7137 | 0.3944 | ||||
+| EDSR [9] | 34.31 | 0.9337 | 0.0951 | 29.98 | 0.8464 | 0.2048 | 27.70 | 0.7644 | 0.3002 | ||||
+| SRGAN [8] | 34.07 | 0.9317 | 0.1007 | 29.79 | 0.8427 | 0.2207 | 27.53 | 0.7571 | 0.3181 | ||||
+| SwinIR [12] | 34.35 | 0.9344 | 0.0942 | 30.06 | 0.8514 | 0.2018 | 27.75 | 0.7684 | 0.2956 | ||||
+| HAT [13] | 34.30 | 0.9339 | 0.0936 | 30.08 | 0.8527 | 0.1979 | 27.77 | 0.7700 | 0.2914 | ||||
+| HAUNet [25] | 33.68 | 0.9270 | 0.1113 | 29.47 | 0.8343 | 0.2333 | 27.25 | 0.7479 | 0.3381 | ||||
+| SPT [27] | 34.12 | 0.9319 | 0.0968 | 29.79 | 0.8439 | 0.2115 | 27.55 | 0.7613 | 0.3072 | ||||
+| ConvFormerSR [28] | 34.27 | 0.9330 | 0.0945 | 29.89 | 0.8461 | 0.2123 | 27.59 | 0.7614 | 0.3109 | ||||
+| MambaIR [15] | 34.52 | 0.9356 | 0.0898 | 30.14 | 0.8513 | 0.1934 | 27.75 | 0.7694 | 0.2874 | ||||
+| FreMamba [16] | 33.87 | 0.9283 | 0.1053 | 29.95 | 0.8471 | 0.2025 | 27.61 | 0.7639 | 0.2992 | ||||
+| GDSR (Ours) | 34.53 | 0.9354 | 0.0918 | 30.15 | 0.8500 | 0.2057 | 27.79 | 0.7683 | 0.3099 | ||||
+| Bicubic |
+22.46 | 0.5128 | 0.6323 | 21.35 | 0.4678 | 0.7330 | 20.86 | 0.4558 | 0.7846 | |||
+| SRCNN [6] | 22.71 | 0.5392 | 0.6277 | 21.48 | 0.4860 | 0.7077 | 20.96 | 0.4702 | 0.7593 | ||||
+| EDSR [9] | 22.99 | 0.5635 | 0.5485 | 21.67 | 0.5056 | 0.6402 | 21.10 | 0.4850 | 0.6917 | ||||
+| SRGAN [8] | 22.96 | 0.5603 | 0.5675 | 21.65 | 0.5041 | 0.6569 | 21.10 | 0.4836 | 0.7051 | ||||
+| SwinIR [12] | 23.09 | 0.5682 | 0.5496 | 21.77 | 0.5096 | 0.6481 | 21.18 | 0.4882 | 0.6900 | ||||
+| HAT [13] | 23.22 | 0.5758 | 0.5338 | 21.82 | 0.5129 | 0.6374 | 21.18 | 0.4895 | 0.6851 | ||||
+| HAUNet [25] | 23.03 | 0.5620 | 0.5635 | 21.69 | 0.5033 | 0.6655 | 21.07 | 0.4824 | 0.7117 | ||||
+| SPT [27] | 23.22 | 0.5724 | 0.5464 | 21.82 | 0.5083 | 0.6565 | 21.26 | 0.4884 | 0.7055 | ||||
+| ConvFormerSR [28] | 23.30 | 0.5739 | 0.5453 | 21.91 | 0.5124 | 0.6459 | 21.27 | 0.4898 | 0.6951 | ||||
+| MambaIR [15] | 23.23 | 0.5778 | 0.5254 | 21.76 | 0.5129 | 0.6247 | 21.16 | 0.4903 | 0.6749 | ||||
+| FreMamba [16] | 23.41 | 0.5783 | 0.5287 | 21.73 | 0.5129 | 0.6200 | 21.12 | 0.4894 | 0.6723 | ||||
+| GDSR (Ours) | 23.45 | 0.5836 | 0.5304 | 21.93 | 0.5156 | 0.6323 | 21.23 | 0.4894 | 0.6828 | ||||
+| Methods | Type | RSSRD-QH x2 | RSSRD-QH x3 | RSSRD-QH x4 | |||||||||
+| PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |||||
+| Bicubic |
+29.37 | 0.8149 | 0.2564 | 26.45 | 0.6447 | 0.4243 | 24.92 | 0.5224 | 0.5764 | |||
+| SRCNN [6] | 29.58 | 0.8218 | 0.3028 | 26.54 | 0.6512 | 0.4241 | 24.98 | 0.5268 | 0.5613 | ||||
+| EDSR [9] | 29.83 | 0.8299 | 0.2838 | 26.70 | 0.6598 | 0.4141 | 25.10 | 0.5343 | 0.5487 | ||||
+| SRGAN [8] | 29.80 | 0.8293 | 0.2859 | 26.68 | 0.6586 | 0.4170 | 25.08 | 0.5333 | 0.5522 | ||||
+| SwinIR [12] | 29.83 | 0.8305 | 0.2844 | 26.70 | 0.6597 | 0.4133 | 25.10 | 0.5349 | 0.5474 | ||||
+| HAT [13] | 29.84 | 0.8307 | 0.2836 | 26.71 | 0.6602 | 0.4111 | 25.10 | 0.5352 | 0.5454 | ||||
+| HAUNet [25] | 29.79 | 0.8289 | 0.2888 | 26.65 | 0.6569 | 0.4177 | 25.07 | 0.5326 | 0.5494 | ||||
+| SPT [27] | 29.83 | 0.8306 | 0.2875 | 26.69 | 0.6602 | 0.4161 | 25.10 | 0.5349 | 0.5478 | ||||
+| ConvFormerSR [28] | 29.82 | 0.8310 | 0.2786 | 26.69 | 0.6603 | 0.4090 | 25.10 | 0.5354 | 0.5452 | ||||
+| MambaIR [15] | 29.86 | 0.8311 | 0.2825 | 26.72 | 0.6607 | 0.4124 | 25.11 | 0.5359 | 0.5445 | ||||
+| FreMamba [16] | 29.80 | 0.8300 | 0.2811 | 26.70 | 0.6601 | 0.4112 | 25.11 | 0.5366 | 0.5414 | ||||
+| GDSR (Ours) | 29.87 | 0.8310 | 0.2823 | 26.73 | 0.6601 | 0.4130 | 25.12 | 0.5354 | 0.5465 | ||||
+| Bicubic |
+23.10 | 0.3991 | 0.5996 | 22.34 | 0.3525 | 0.6993 | 21.85 | 0.3284 | 0.7764 | |||
+| SRCNN [6] | 23.27 | 0.4159 | 0.6520 | 22.44 | 0.3588 | 0.7523 | 21.91 | 0.3311 | 0.8078 | ||||
+| EDSR [9] | 23.67 | 0.4485 | 0.6269 | 22.74 | 0.3817 | 0.7283 | 22.15 | 0.3444 | 0.8028 | ||||
+| SRGAN [8] | 23.57 | 0.4387 | 0.6442 | 22.64 | 0.3759 | 0.7421 | 22.09 | 0.3408 | 0.8174 | ||||
+| SwinIR [12] | 23.68 | 0.4463 | 0.6232 | 22.72 | 0.3780 | 0.7432 | 22.15 | 0.3440 | 0.8104 | ||||
+| HAT [13] | 23.75 | 0.4528 | 0.6120 | 22.79 | 0.3823 | 0.7312 | 22.18 | 0.3456 | 0.8061 | ||||
+| HAUNet [25] | 23.72 | 0.4501 | 0.6136 | 22.76 | 0.3797 | 0.7360 | 22.16 | 0.3442 | 0.8096 | ||||
+| SPT [27] | 23.58 | 0.4353 | 0.6512 | 22.73 | 0.3743 | 0.7520 | 22.14 | 0.3415 | 0.8144 | ||||
+| ConvFormerSR [28] | 23.67 | 0.4398 | 0.6267 | 22.77 | 0.3779 | 0.7365 | 22.18 | 0.3436 | 0.8051 | ||||
+| MambaIR [15] | 23.82 | 0.4568 | 0.6036 | 22.84 | 0.3835 | 0.7387 | 22.23 | 0.3476 | 0.8003 | ||||
+| FreMamba [16] | 23.59 | 0.4329 | 0.6009 | 22.60 | 0.3660 | 0.7242 | 22.06 | 0.3358 | 0.7894 | ||||
+| GDSR (Ours) | 23.75 | 0.4454 | 0.6197 | 22.87 | 0.3864 | 0.7308 | 22.27 | 0.3488 | 0.8044 | ||||
 | Categoties | Bicubic | SRGAN [8] | EDSR [9] | HAT [13] | MambaIR [15] | ConvFormerSR [28] | FreMamba [16] | GDSR (Ours) | ||||||||
 | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |
-Airport | 27.43 | 0.7387 | 28.48 | 0.7765 | 28.63 | 0.7823 | 28.69 | 0.7837 | 28.72 | 0.7856 | 28.59 | 0.7805 | 28.68 | 0.7851 | 28.72 | 0.7853 |
-Bare Land | 34.45 | 0.8344 | 34.84 | 0.8422 | 34.92 | 0.8439 | 34.94 | 0.8445 | 34.96 | 0.8449 | 34.91 | 0.8436 | 34.93 | 0.8444 | 34.97 | 0.8449 |
-Baseball Field | 29.19 | 0.7786 | 30.27 | 0.8094 | 30.42 | 0.8137 | 30.45 | 0.8144 | 30.52 | 0.8160 | 30.35 | 0.8123 | 30.43 | 0.8147 | 30.53 | 0.8160 |
-Beach | 31.32 | 0.7947 | 31.86 | 0.8082 | 31.94 | 0.8103 | 31.99 | 0.8117 | 32.00 | 0.8118 | 31.94 | 0.8096 | 31.95 | 0.8113 | 32.00 | 0.8115 |
-Bridge | 30.19 | 0.8063 | 31.27 | 0.8312 | 31.46 | 0.8353 | 31.52 | 0.8363 | 31.56 | 0.8372 | 31.42 | 0.8339 | 31.52 | 0.8368 | 31.57 | 0.8371 |
-Center | 25.76 | 0.6980 | 27.18 | 0.7590 | 27.38 | 0.7665 | 27.47 | 0.7694 | 27.51 | 0.7706 | 27.33 | 0.7645 | 27.43 | 0.7692 | 27.51 | 0.7705 |
-Church | 23.09 | 0.6084 | 24.41 | 0.6807 | 24.56 | 0.6895 | 24.61 | 0.6920 | 24.65 | 0.6941 | 24.52 | 0.6875 | 24.56 | 0.6919 | 24.67 | 0.6947 |
-Commercial | 26.72 | 0.7178 | 27.59 | 0.7582 | 27.73 | 0.7645 | 27.77 | 0.7660 | 27.81 | 0.7679 | 27.68 | 0.7623 | 27.72 | 0.7659 | 27.81 | 0.7681 |
-D-Residential | 23.45 | 0.6301 | 24.46 | 0.6922 | 24.60 | 0.7004 | 24.64 | 0.7026 | 24.69 | 0.7051 | 24.56 | 0.6985 | 24.60 | 0.7029 | 24.70 | 0.7056 |
-Desert | 37.14 | 0.8869 | 37.54 | 0.8933 | 37.58 | 0.8944 | 37.63 | 0.8951 | 37.62 | 0.8949 | 37.58 | 0.8941 | 37.60 | 0.8946 | 37.62 | 0.8950 |
-Farmland | 32.32 | 0.7934 | 33.18 | 0.8166 | 33.33 | 0.8212 | 33.35 | 0.8216 | 33.42 | 0.8236 | 33.30 | 0.8202 | 33.39 | 0.8235 | 33.43 | 0.8240 |
-Forest | 27.67 | 0.6411 | 27.84 | 0.6514 | 27.90 | 0.6558 | 27.95 | 0.6605 | 27.95 | 0.6604 | 27.88 | 0.6552 | 27.92 | 0.6598 | 27.97 | 0.6616 |
-Industrial | 26.29 | 0.7063 | 27.43 | 0.7566 | 27.62 | 0.7647 | 27.67 | 0.7669 | 27.71 | 0.7689 | 27.56 | 0.7621 | 27.62 | 0.7665 | 27.72 | 0.7690 |
-Meadow | 32.07 | 0.7100 | 32.33 | 0.7136 | 32.38 | 0.7157 | 32.40 | 0.7167 | 32.41 | 0.7172 | 32.37 | 0.7154 | 32.39 | 0.7175 | 32.42 | 0.7172 |
-M-Residential | 26.32 | 0.6578 | 27.40 | 0.7055 | 27.55 | 0.7124 | 27.58 | 0.7136 | 27.63 | 0.7159 | 27.52 | 0.7113 | 27.55 | 0.7140 | 27.64 | 0.7164 |
-Mountain | 28.05 | 0.6978 | 28.41 | 0.7127 | 28.48 | 0.7163 | 28.50 | 0.7184 | 28.50 | 0.7183 | 28.46 | 0.7152 | 28.45 | 0.7171 | 28.50 | 0.7183 |
-Park | 27.14 | 0.7060 | 27.77 | 0.7332 | 27.87 | 0.7383 | 27.91 | 0.7406 | 27.93 | 0.7415 | 27.83 | 0.7365 | 27.86 | 0.7395 | 27.94 | 0.7422 |
-Parking | 24.11 | 0.7310 | 26.18 | 0.8059 | 26.51 | 0.8159 | 26.68 | 0.8205 | 26.76 | 0.8225 | 26.46 | 0.8145 | 26.70 | 0.8226 | 26.77 | 0.8228 |
-Playground | 29.33 | 0.7723 | 30.70 | 0.8132 | 30.94 | 0.8200 | 31.03 | 0.8216 | 31.10 | 0.8235 | 30.88 | 0.8182 | 31.03 | 0.8230 | 31.12 | 0.8237 |
-Pond | 28.76 | 0.7559 | 29.38 | 0.7749 | 29.46 | 0.7781 | 29.49 | 0.7790 | 29.51 | 0.7796 | 29.44 | 0.7769 | 29.46 | 0.7788 | 29.51 | 0.7796 |
-Port | 25.83 | 0.7727 | 26.94 | 0.8207 | 27.11 | 0.8255 | 27.18 | 0.8273 | 27.21 | 0.8283 | 27.05 | 0.8237 | 27.13 | 0.8267 | 27.21 | 0.8278 |
-Railway Station | 27.00 | 0.6965 | 27.87 | 0.7347 | 28.04 | 0.7434 | 28.11 | 0.7456 | 28.13 | 0.7474 | 27.99 | 0.7404 | 28.07 | 0.7461 | 28.14 | 0.7477 |
-Resort | 26.78 | 0.7160 | 27.67 | 0.7547 | 27.81 | 0.7601 | 27.86 | 0.7622 | 27.89 | 0.7634 | 27.78 | 0.7588 | 27.82 | 0.7619 | 27.91 | 0.7638 |
-River | 29.29 | 0.7201 | 29.74 | 0.7378 | 29.81 | 0.7412 | 29.84 | 0.7426 | 29.85 | 0.7430 | 29.79 | 0.7401 | 29.80 | 0.7422 | 29.86 | 0.7431 |
-School | 25.81 | 0.7046 | 26.81 | 0.7503 | 26.96 | 0.7568 | 27.00 | 0.7587 | 27.04 | 0.7604 | 26.91 | 0.7546 | 26.94 | 0.7577 | 27.05 | 0.7607 |
-S-Residential | 25.71 | 0.5790 | 26.15 | 0.6001 | 26.22 | 0.6055 | 26.25 | 0.6070 | 26.27 | 0.6079 | 26.21 | 0.6046 | 26.23 | 0.6075 | 26.27 | 0.6077 |
-Square | 27.51 | 0.7286 | 28.71 | 0.7782 | 28.90 | 0.7847 | 28.96 | 0.7870 | 28.99 | 0.7881 | 28.85 | 0.7831 | 28.92 | 0.7868 | 29.01 | 0.7882 |
-Stadium | 26.23 | 0.7241 | 27.55 | 0.7771 | 27.75 | 0.7845 | 27.82 | 0.7873 | 27.85 | 0.7882 | 27.68 | 0.7823 | 27.78 | 0.7867 | 27.85 | 0.7881 |
-Storage Tanks | 25.15 | 0.6667 | 26.25 | 0.7190 | 26.39 | 0.7255 | 26.41 | 0.7264 | 26.45 | 0.7287 | 26.33 | 0.7233 | 26.39 | 0.7274 | 26.45 | 0.7286 |
-Viaduct | 26.77 | 0.6817 | 27.70 | 0.7225 | 27.88 | 0.7310 | 27.91 | 0.7325 | 27.98 | 0.7361 | 27.81 | 0.7277 | 27.87 | 0.7326 | 27.98 | 0.7363 |
-Average | 27.89 | 0.7238 | 28.80 | 0.7597 | 28.94 | 0.7653 | 28.99 | 0.7671 | 29.02 | 0.7684 | 28.90 | 0.7637 | 28.96 | 0.7672 | 29.03 | 0.7686 |
-IV-C2 Effect of Wavelet Loss
-To evaluate the effectiveness of incorporating Wavelet Loss in training, ablation experiments were conducted comparing models trained with Wavelet Loss to those using only Loss.
-These experiments spanned multiple datasets and degradation models, employing a variety of architectures.
-The results consistently demonstrated that Wavelet Loss improves PSNR, with SSIM also showing notable enhancement in most cases.
-On the RSSRD-QH dataset, a decrease in SSIM performance was observed, which was attributed to the complexity of basin images, characterized by intricate textures and diverse structures.
-Wavelet Loss emphasized high-frequency details, which didn’t align with SSIM’s focus on global structural similarity.
-A decline in LPIPS was also observed across all experiments with Wavelet Loss, due to its emphasis on high-frequency details and structural accuracy.
-The prioritization of fine-grained features introduced perceptual differences, as captured by LPIPS.
-Nonetheless, the substantial improvements in PSNR and SSIM underscored the overall effectiveness of Wavelet Loss in producing high-quality SR images, guiding the correct reconstruction of textures across various architectures.
-As shown in Figure 9, despite MambaIR’s global receptive field and its superior performance in experiments, the model still struggled to accurately reconstruct certain textures without Wavelet Loss.
-The inclusion of Wavelet Loss significantly improved the reconstruction, particularly in the high-frequency components (LH, HL, HH), where fine-grained details were preserved.
-This highlights Wavelet Loss’s role in guiding the model to recover intricate structures and accurately reconstruct complex textures, thereby enhancing overall image quality.
+| Airport | 27.43 | 0.7387 | 28.48 | 0.7765 | 28.63 | 0.7823 | 28.69 | 0.7837 | 28.72 | 0.7856 | 28.59 | 0.7805 | 28.68 | 0.7851 | 28.74 | 0.7865 |
+| Bare Land | 34.45 | 0.8344 | 34.84 | 0.8422 | 34.92 | 0.8439 | 34.94 | 0.8445 | 34.96 | 0.8449 | 34.91 | 0.8436 | 34.93 | 0.8444 | 34.98 | 0.8455 |
+| Baseball Field | 29.19 | 0.7786 | 30.27 | 0.8094 | 30.42 | 0.8137 | 30.45 | 0.8144 | 30.52 | 0.8160 | 30.35 | 0.8123 | 30.43 | 0.8147 | 30.55 | 0.8169 |
+| Beach | 31.32 | 0.7947 | 31.86 | 0.8082 | 31.94 | 0.8103 | 31.99 | 0.8117 | 32.00 | 0.8118 | 31.94 | 0.8096 | 31.95 | 0.8113 | 32.03 | 0.8124 |
+| Bridge | 30.19 | 0.8063 | 31.27 | 0.8312 | 31.46 | 0.8353 | 31.52 | 0.8363 | 31.56 | 0.8372 | 31.42 | 0.8339 | 31.52 | 0.8368 | 31.59 | 0.8378 |
+| Center | 25.76 | 0.6980 | 27.18 | 0.7590 | 27.38 | 0.7665 | 27.47 | 0.7694 | 27.51 | 0.7706 | 27.33 | 0.7645 | 27.43 | 0.7692 | 27.52 | 0.7714 |
+| Church | 23.09 | 0.6084 | 24.41 | 0.6807 | 24.56 | 0.6895 | 24.61 | 0.6920 | 24.65 | 0.6941 | 24.52 | 0.6875 | 24.56 | 0.6919 | 24.68 | 0.6954 |
+| Commercial | 26.72 | 0.7178 | 27.59 | 0.7582 | 27.73 | 0.7645 | 27.77 | 0.7660 | 27.81 | 0.7679 | 27.68 | 0.7623 | 27.72 | 0.7659 | 27.83 | 0.7692 |
+| D-Residential | 23.45 | 0.6301 | 24.46 | 0.6922 | 24.60 | 0.7004 | 24.64 | 0.7026 | 24.69 | 0.7051 | 24.56 | 0.6985 | 24.60 | 0.7029 | 24.72 | 0.7067 |
+| Desert | 37.14 | 0.8869 | 37.54 | 0.8933 | 37.58 | 0.8944 | 37.63 | 0.8951 | 37.62 | 0.8949 | 37.58 | 0.8941 | 37.60 | 0.8946 | 37.65 | 0.8954 |
+| Farmland | 32.32 | 0.7934 | 33.18 | 0.8166 | 33.33 | 0.8212 | 33.35 | 0.8216 | 33.42 | 0.8236 | 33.30 | 0.8202 | 33.39 | 0.8235 | 33.45 | 0.8245 |
+| Forest | 27.67 | 0.6411 | 27.84 | 0.6514 | 27.90 | 0.6558 | 27.95 | 0.6605 | 27.95 | 0.6604 | 27.88 | 0.6552 | 27.92 | 0.6598 | 27.99 | 0.6617 |
+| Industrial | 26.29 | 0.7063 | 27.43 | 0.7566 | 27.62 | 0.7647 | 27.67 | 0.7669 | 27.71 | 0.7689 | 27.56 | 0.7621 | 27.62 | 0.7665 | 27.74 | 0.7701 |
+| Meadow | 32.07 | 0.7100 | 32.33 | 0.7136 | 32.38 | 0.7157 | 32.40 | 0.7167 | 32.41 | 0.7172 | 32.37 | 0.7154 | 32.39 | 0.7175 | 32.43 | 0.7176 |
+| M-Residential | 26.32 | 0.6578 | 27.40 | 0.7055 | 27.55 | 0.7124 | 27.58 | 0.7136 | 27.63 | 0.7159 | 27.52 | 0.7113 | 27.55 | 0.7140 | 27.66 | 0.7170 |
+| Mountain | 28.05 | 0.6978 | 28.41 | 0.7127 | 28.48 | 0.7163 | 28.50 | 0.7184 | 28.50 | 0.7183 | 28.46 | 0.7152 | 28.45 | 0.7171 | 28.52 | 0.7192 |
+| Park | 27.14 | 0.7060 | 27.77 | 0.7332 | 27.87 | 0.7383 | 27.91 | 0.7406 | 27.93 | 0.7415 | 27.83 | 0.7365 | 27.86 | 0.7395 | 27.96 | 0.7431 |
+| Parking | 24.11 | 0.7310 | 26.18 | 0.8059 | 26.51 | 0.8159 | 26.68 | 0.8205 | 26.76 | 0.8225 | 26.46 | 0.8145 | 26.70 | 0.8226 | 26.81 | 0.8238 |
+| Playground | 29.33 | 0.7723 | 30.70 | 0.8132 | 30.94 | 0.8200 | 31.03 | 0.8216 | 31.10 | 0.8235 | 30.88 | 0.8182 | 31.03 | 0.8230 | 31.13 | 0.8244 |
+| Pond | 28.76 | 0.7559 | 29.38 | 0.7749 | 29.46 | 0.7781 | 29.49 | 0.7790 | 29.51 | 0.7796 | 29.44 | 0.7769 | 29.46 | 0.7788 | 29.53 | 0.7802 |
+| Port | 25.83 | 0.7727 | 26.94 | 0.8207 | 27.11 | 0.8255 | 27.18 | 0.8273 | 27.21 | 0.8283 | 27.05 | 0.8237 | 27.13 | 0.8267 | 27.22 | 0.8280 |
+| Railway Station | 27.00 | 0.6965 | 27.87 | 0.7347 | 28.04 | 0.7434 | 28.11 | 0.7456 | 28.13 | 0.7474 | 27.99 | 0.7404 | 28.07 | 0.7461 | 28.16 | 0.7489 |
+| Resort | 26.78 | 0.7160 | 27.67 | 0.7547 | 27.81 | 0.7601 | 27.86 | 0.7622 | 27.89 | 0.7634 | 27.78 | 0.7588 | 27.82 | 0.7619 | 27.92 | 0.7644 |
+| River | 29.29 | 0.7201 | 29.74 | 0.7378 | 29.81 | 0.7412 | 29.84 | 0.7426 | 29.85 | 0.7430 | 29.79 | 0.7401 | 29.80 | 0.7422 | 29.88 | 0.7439 |
+| School | 25.81 | 0.7046 | 26.81 | 0.7503 | 26.96 | 0.7568 | 27.00 | 0.7587 | 27.04 | 0.7604 | 26.91 | 0.7546 | 26.94 | 0.7577 | 27.07 | 0.7619 |
+| S-Residential | 25.71 | 0.5790 | 26.15 | 0.6001 | 26.22 | 0.6055 | 26.25 | 0.6070 | 26.27 | 0.6079 | 26.21 | 0.6046 | 26.23 | 0.6075 | 26.29 | 0.6080 |
+| Square | 27.51 | 0.7286 | 28.71 | 0.7782 | 28.90 | 0.7847 | 28.96 | 0.7870 | 28.99 | 0.7881 | 28.85 | 0.7831 | 28.92 | 0.7868 | 29.03 | 0.7891 |
+| Stadium | 26.23 | 0.7241 | 27.55 | 0.7771 | 27.75 | 0.7845 | 27.82 | 0.7873 | 27.85 | 0.7882 | 27.68 | 0.7823 | 27.78 | 0.7867 | 27.87 | 0.7889 |
+| Storage Tanks | 25.15 | 0.6667 | 26.25 | 0.7190 | 26.39 | 0.7255 | 26.41 | 0.7264 | 26.45 | 0.7287 | 26.33 | 0.7233 | 26.39 | 0.7274 | 26.47 | 0.7298 |
+| Viaduct | 26.77 | 0.6817 | 27.70 | 0.7225 | 27.88 | 0.7310 | 27.91 | 0.7325 | 27.98 | 0.7361 | 27.81 | 0.7277 | 27.87 | 0.7326 | 28.00 | 0.7378 |
+| Average | 27.89 | 0.7238 | 28.80 | 0.7597 | 28.94 | 0.7653 | 28.99 | 0.7671 | 29.02 | 0.7684 | 28.90 | 0.7637 | 28.96 | 0.7672 | 29.05 | 0.7694 |
 IV-D Comparisons With State-of-the-Art
 IV-D1 Comparative Methods
-CNN-based models, including SRCNN [6] and EDSR [9]; GAN-based models, including SRGAN [8]; Transformer-based models, including SwinIR [14] and HAT [15]; CNN-Transformer hybrid models, including ConvFormerSR [16]; and Mamba-based models, including MambaIR [18] and FreMamba [20], were involved in evaluating the SR performance of our GDSR against SOTA methods on RSI-SR.
-We also report the results of our model variants, GDSR_TC and GDSR_MC, as well as the large-scale variant GDSR_L.
+CNN-based models, including SRCNN [6], EDSR [9], and HAUNet [25]; GAN-based models, including SRGAN [8]; Transformer-based models, including SwinIR [12], HAT [13], and SPT [27]; CNN-Transformer hybrid models, including ConvFormerSR [28]; and Mamba-based models, including MambaIR [15] and FreMamba [16], were involved in evaluating the SR performance of our GDSR against SOTA methods on RSI-SR.
 IV-D2 Quantitative Evaluations
-The quantitative results on the AID and RSSRD-QH datasets are presented in Table V.
-Our proposed GDSR and GDSR_L models achieved superior performance across all metrics, demonstrating their remarkable SR capabilities in diverse remote sensing tests.
-On the AID and RSSRD-QH datasets, GDSR achieved performance comparable to MambaIR.
-On the AID_CDM and RSSRD-QH_CDM datasets, GDSR’s advantages became more pronounced.
-It achieved PSNR, SSIM, and LPIPS values of 24.10, 0.5770, and 0.5986, respectively, on AID_CDM, and PSNR and SSIM values of 22.86 and 0.3861, respectively, on RSSRD-QH_CDM, significantly outperforming other methods.
-GDSR_L consistently led across all tests, and both GDSR_TC and GDSR_MC achieved performance comparable to FreMamba on the AID and RSSRD-QH datasets while demonstrating substantial improvements on the AID_CDM and RSSRD-QH_CDM datasets.
-Specifically, GDSR_MC achieved a 0.14 dB PSNR improvement on AID_CDM and a 0.24 dB improvement on RSSRD-QH_CDM compared to FreMamba.
-To further evaluate the generalization capability of the super-resolution models across diverse remote sensing scenarios, we report the PSNR and SSIM results across 30 scene categories in the AID dataset, as shown in Table VII.
+The quantitative evaluation results across different datasets and scaling factors were comprehensively analyzed through Tables VII and VIII to assess reconstruction accuracy and perceptual quality.
+On the AID dataset under bicubic degradation, the proposed GDSR demonstrated state-of-the-art performance with PSNR values of 31.81 dB (×2), 29.05 dB (×3), and 27.39 dB (×4), showing improvements of 0.04-0.11 dB over transformer-based HAT and CNN-based EDSR.
+The structural preservation capability was further verified by SSIM metrics, where GDSR achieved 0.8592 (×2), 0.7694 (×3), and 0.7022 (×4).
+Although FreMamba exhibited superior LPIPS scores across ×2-4 scales, GDSR maintained competitive perceptual quality while balancing pixel-wise accuracy and visual realism.
+For more challenging CDM-synthesized data, GDSR showed remarkable robustness with consistent performance gains, surpassing MambaIR by 0.11 dB PSNR and 0.26% SSIM in ×2 super-resolution.
+The superiority became more evident on the UCMerced dataset, where GDSR established new benchmarks in 7 out of 12 metrics.
+Notably, it achieved 0.17 dB PSNR improvement over MambaIR and 0.11 dB gain against HAT in ×3 super-resolution under CDM degradation.
+Evaluation on the real-world RSSRD-QH dataset further validated the practical effectiveness of GDSR.
+Under bicubic degradation, the method attained peak PSNR values of 29.87 dB (×2), 26.73 dB (×3), and 25.12 dB (×4).
+In CDM degradation scenarios, GDSR exhibited particularly strong performance at high super-resolution factors, reaching 22.87 dB PSNR (×3) and 22.27 dB PSNR (×4) with corresponding SSIM values of 0.3864 and 0.3488.
+To further evaluate the generalization capability of the super-resolution models across diverse remote sensing scenarios, we report the PSNR and SSIM results across 30 scene categories in the AID dataset under bicubic degradation, as shown in Table IX.
 As evidenced, GDSR exhibited stronger generalization capability compared to state-of-the-art methods, achieving the best performance in nearly all remote sensing scenarios.
-Specifically, for the ”Parking” and ”Playground” scenes, GDSR outperformed HAT by 0.09 dB in PSNR and by 0.0023 and 0.0021 in SSIM, respectively.
+Specifically, for the ”Parking” and ”Playground” scenes, GDSR outperformed HAT by 0.13 dB and 0.1 dB in PSNR and by 0.33% and 0.28% in SSIM, respectively.
 Additionally, in the most complex categories, such as ”Church” and ”DenseResidential,” GDSR also achieved the best performance.
@@
 This aligned with our experimental results on the RGEG Branch and RDEG Branch, where larger receptive fields enabled the activation of more pixels for image reconstruction, enhancing performance.
-However, the results from MambaIR and FreMamba, along with the experiments on F2B and F2C, indicated that spatial global modeling in complex remote sensing images had reached a bottleneck.
+However, the results from MambaIR and FreMamba, along with the experiments on F2B and F2C, indicated that spatial global modeling in complex RSIs had reached a bottleneck.
 Simply increasing the receptive field no longer yielded significant performance gains.
 A more practical solution was to effectively utilize large receptive fields to activate the appropriate regions and pixels for accurate image reconstruction.
-The outstanding generalization capability and superior performance of GDSR aligned with our motivation to address the challenges of large-scale remote sensing images efficiently through the introduction of a global-detail dual-branch structure.
+The outstanding generalization capability and superior performance of GDSR align with our motivation to address the challenges of large-scale RSIs efficiently through the introduction of a global-detail dual-branch structure.
 IV-D3 Complexity and Efficiency Evaluation
 To evaluate the complexity and computational efficiency of our proposed models, we conducted quantitative assessments across various metrics.
-As shown in Table VIII, the bicubic interpolation technique inherently avoids the need for parameter optimization, resulting in the fastest inference speed within our evaluation scope.
+As shown in Table X, the bicubic interpolation technique inherently avoids the need for parameter optimization, resulting in the fastest inference speed within our evaluation scope.
 The CNN-based model achieved the highest inference speed among all tested deep learning models, primarily due to the inherent efficiency of convolutional parallel operations.
 The EDSR model, with its deep residual structure, had the largest number of parameters and FLOPs, yet achieved inference speeds several times faster than other architectures.
+| Model | #Param. | FLOPs | Memory | FPS |
+| Bicubic | - | - | 22MB | 27498.0 |
+| SRCNN [6] | 0.02M | 4.65G | 140MB | 1109.0 |
+| SRGAN [8] | 1.55M | 47.86G | 300MB | 323.3 |
+| EDSR [9] | 43.68M | 1119.98G | 1068MB | 35.1 |
+| SwinIR [12] | 16.62M | 446.67G | 516MB | 5.5 |
+| HAT [13] | 20.81M | 652.17G | 1446MB | 4.6 |
+| HAUNet [25] | 2.64M | 1353.07G | 7658MB | 3.1 |
+| SPT [27] | 3.25M | 668.46G | 9974MB | 2.2 |
+| ConvFormerSR [28] | 16.49M | 442.18G | 500MB | 7.0 |
+| MambaIR [15] | 15.11M | 372.85G | 556MB | 6.4 |
+| FreMamba [16] | 11.46M | 229.50G | 546MB | 5.5 |
+| GDSR (Ours) | 13.17M | 338.73G | 518MB | 14.8 |
 Transformer-based models, such as SwinIR and HAT, exhibited slower inference speeds compared to CNN-based models but achieved significant performance gains.
@@
 GDSR, by incorporating RWKV with CNN, not only reduced the parameter count but also achieved inference speeds several times faster than Transformer- and Mamba-based models, while delivering enhanced performance.
-GDSR_TC and GDSR_MC further optimized performance by combining Swin Transformer and Mamba with CNN, leading to a reduction in parameter size and improved inference speed, without sacrificing performance.
-GDSR_L, which increased the network depth, consistently outperformed other models in experimental evaluations while maintaining reasonable inference speeds.
-As demonstrated in Tables II, III, and IV, these results underscore the superiority of the Global-Detail dual-branch architecture design.
-Comparing Tables VIII and V reveals that the complexity of our proposed model is comparable to most state-of-the-art models, while its efficiency is superior, and its performance is significantly better.
+As demonstrated in Tables II, III, and V, these results underscore the superiority of the Global-Detail dual-branch architecture design.
+Comparing Tables X and VII reveals that the complexity of our proposed model is comparable to most state-of-the-art models, while its efficiency is superior, and its performance is significantly better.
 These results highlight the practical potential of our model, demonstrating an effective balance between computational efficiency and performance, making it well-suited for real-world applications.
-Model | #Param. | FLOPs | FPS |
-Bicubic | - | - | 21505.6 |
-SRCNN [6] | 0.02M | 10.47G | 533.1 |
-SRGAN [8] | 1.55M | 107.68G | 151.9 |
-EDSR [9] | 43.68M | 2519.95G | 15.5 |
-SwinIR [14] | 16.62M | 1005.01G | 2.3 |
-HAT [15] | 20.81M | 1467.38G | 2.2 |
-MambaIR [18] | 15.11M | 838.91G | 2.8 |
-ConvFormerSR [16] | 16.49M | 994.90G | 3.2 |
-FreMamba [20] | 11.46M | 516.21G | 2.5 |
-GDSR_TC (Ours) | 12.83M | 758.82G | 3.3 |
-GDSR_MC (Ours) | 13.99M | 788.28G | 4.1 |
-GDSR (Ours) | 13.17M | 762.14G | 6.4 |
-GDSR_L (Ours) | 25.86M | 1496.14G | 3.2 |
 IV-D4 Qualitative Results
@@
 13, based on Figs.
-10 and 12, demonstrated that GDSR benefitted from the superior long-range modeling capability of RGEG, enabling it to leverage more pixels during the SR process.
+10 and 12, demonstrated that GDSR benefited from the superior long-range modeling capability of RGEG, enabling it to leverage more pixels during the SR process.
 Additionally, the strong detail extraction capability of RDEG allowed GDSR to utilize pixels correctly for reconstruction.
-Specifically, for the RSSRD-QH_CDM dataset, when large-scale and global information was present, CNN- and Transformer-based SR networks struggled to extract sufficient information, and while the Mamba-based SR network activated a large number of pixels, they failed to use them correctly for reconstruction.
+Specifically, for the RSSRD-QH_CDM dataset, when large-scale and global information was present, CNN- and Transformer-based SR networks struggled to extract sufficient information, and while the Mamba-based SR network activated a large number of pixels, it failed to use them correctly for reconstruction.
 In contrast, thanks to the effective integration of RGEG and RDEG, the GDSR-generated results significantly outperformed other SR models.
@@
 IV-D5 Real-World Image Testing on the Sanjiangyuan Region
-To further evaluate the model’s performance in real-world scenarios, we tested it on remote sensing images from real-world environments.
+To further evaluate the model’s performance in real-world scenarios, we tested it on RSIs from real-world environments.
 As shown in Fig.
@@
 This improvement is crucial for accurate environmental analysis, as the restored landscape provides more precise data for studying biodiversity conservation and the enhancement of ecological functions in the alpine meadow ecosystem of Qinghai Province.
-V Conclusion
-In this study, we first introduce the Receptance Weighted Key Value for remote sensing image super-resolution.
-GDSR effectively models the global dependencies of large-scale remote sensing images while incorporating local details, all with linear complexity and significantly improved inference speed.
-Specifically, we design an efficient and effective Global-Detail dual-branch architecture that correctly activates more pixels spatially for image reconstruction.
+V Discussion
+The proposed method demonstrated consistent SR performance across diverse remote sensing scenarios, validated on the AID (30 land-use categories) [44], UCMerced (21-class geospatial scenes) [45], and custom RSSRD-QH (Qinghai Plateau ecology) datasets.
+By enhancing texture fidelity in complex environments such as fragmented agricultural landscapes, high-density urban settlements, and ecologically sensitive alpine regions, this approach supports critical applications requiring precise feature extraction.
+Improved road network reconstruction in severely degraded AID samples aids emergency routing during disasters, while sharpened landscape details in RSSRD-QH enable accurate environmental monitoring.
+The method’s adaptability to both global benchmarks and region-specific challenges positions it as a practical tool for revitalizing historical mid-resolution archives into high-precision analysis-ready data.
+Additionally, we tested GDSR’s robustness against challenging, remote sensing-specific artifacts.
+While our initial CDM followed established blind SR methods [49, 50, 51, 52], real-world RSIs often contain severe sensor-induced stripe noise.
+To validate our method’s adaptability, we augmented the CDM with this artifact and retrained the GDSR model.
+As demonstrated in Fig.
+15, the results confirmed GDSR’s strong adaptive capabilities.
+The retrained model successfully identified and eliminated severe stripe noise in images, leading to quantitative and visual improvements.
+This positive result paves the way for future work to incorporate other challenging degradations, such as atmospheric scattering, further enhancing the model’s practical utility.
+VI Conclusion
+In this study, we first introduce the RWKV for RSISR.
+GDSR effectively models the global dependencies of large-scale RSI while incorporating local details, all with linear complexity and significantly improved inference speed.
+Specifically, we design an efficient and effective Global-Detail dual-branch architecture that correctly activates more pixels spatially for SR.
 To combine global and local representations, we propose GDRM, which adaptively adjusts the features of the dual branches and fuses them.
-Additionally, we introduce Wavelet Loss to help the model better capture high-frequency details in remote sensing image super-resolution tasks.
-Extensive quantitative and qualitative experiments on the AID, AID_CDM, RSSRD-QH, and RSSRD-QH_CDM benchmarks demonstrate that our GDSR outperforms state-of-the-art CNN-, Transformer-, and Mamba-based SR models in remote sensing image super-resolution tasks.
-The proposed Wavelet Loss also enhances the performance of various models in remote sensing image super-resolution tasks.
-This study focuses solely on ×3 SR, limiting the flexibility to explore SR at different scales.
-Experiments show that excessive high-frequency components introduced by Wavelet Loss can lead to a decline in perceptual quality.
-In future work, we plan to extend our GDSR to include more scaling factors, further demonstrating its robustness and effectiveness.
-Additionally, we aim to optimize Wavelet Loss to improve its perceptual enhancement.
+Additionally, we introduce Dual-Group Multi-Scale Wavelet Loss to help the model better reconstruct high-frequency details.
+Extensive quantitative and qualitative experiments on the AID, UCMerced, and RSSRD-QH benchmark datasets using two degradation methods demonstrate that our GDSR outperforms state-of-the-art CNN-, Transformer-, and Mamba-based SR models in RSI-SR tasks.
+Nevertheless, our GDSR does exhibit some shortcomings.
+First, the use of simple stacked residual convolutions for detail extraction introduces a high parameter count, limiting model efficiency.
+Future work will explore more efficient architectures specifically tailored for RSIs.
+Second, the current validation is limited to RGB RSIs, lacking evaluation on hyperspectral SR scenarios where balancing spatial enhancement with spectral fidelity is paramount.
+Extending the method to hyperspectral data and rigorously assessing its spectral accuracy will be a key future direction.
+Moreover, we consider investigating the model’s applicability to multi-temporal or cross-sensor RSIs, addressing challenges like temporal consistency and domain adaptation to further enhance its practical utility.